### v1

# CLIPArTT: Light-weight Adaptation of CLIP to New Domains at Test Time

Abstract
Pre-trained vision-language models (VLMs), exemplified by CLIP, demonstrate remarkable adaptability across zero-shot classification tasks without additional training. However, their performance diminishes in the presence of domain shifts. In this study, we introduce CLIP Adaptation duRing Test-Time (CLIPArTT), a fully test-time adaptation (TTA) approach for CLIP, which involves automatic text prompts construction during inference for their use as text supervision. Our method employs a unique, minimally invasive text prompt tuning process, wherein multiple predicted classes are aggregated into a single new text prompt, used as pseudo label to re-classify inputs in a transductive manner. Additionally, we pioneer the standardization of TTA benchmarks (e.g., TENT) in the realm of VLMs. Our findings demonstrate that, without requiring additional transformations nor new trainable modules, CLIPArTT enhances performance dynamically across non-corrupted datasets such as CIFAR-10, corrupted datasets like CIFAR-10-C and CIFAR-10.1, alongside synthetic datasets such as VisDA-C. This research underscores the potential for improving VLMs’ adaptability through novel test-time strategies, offering insights for robust performance across varied datasets and environments. The code can be found at: https://github.com/dosowiechi/CLIPArTT.git
1 Introduction
Combining vision and language modalities for learning, namely a Vision Language model (VLM), has demonstrated an outstanding performance in different vision tasks Radford et al. (2021); Jia et al. (2021); Kirillov et al. (2023). Remarkably, these models are surprisingly effective at zero-shot generalization, where a new task can be outside the original scope of the training set, without any additional supervision to fine tune the model. Models such as CLIP Radford et al. (2021) have then been employed in fields as diverse as video recognition Lin et al. (2022), audio Guzhov et al. (2022), and medical imaging Liu et al. (2023). These advancements point these methods to become critical in future machine learning research and applications.
As in other more traditional deep architectures (e.g., CNNs), CLIP is prone to performance degradation on domains to which it was not originally exposed. Recent research trends suggest that domain adaptation mechanisms can play an important role in deploying CLIP Lai et al. (2023); Shu et al. (2022). The challenge, however, is to adapt the model to new domains in real-time and as fast as possible, to maintain its attractive zero-shot capabilities without the need of retraining.
In this paper, CLIP is contextualized in the setting of Test-Time Adaptation, a challenging yet practical scenario of domain adaptation. In realistic scenarios, a model needs to adapt to new data on-the-fly to cope with unknown distribution shifts, and without using any class supervision. Although an experimental boilerplate was standardized in recent years, CLIP has not been integrated to it yet. Additionally, we introduce a powerful adaptation technique that achieves state-of-the-art performance without a significant computational overhead. Comprehensive studies are performed on three datasets containing different types of domain shifts on several levels of severity, resulting in a total of 155 evaluation scenarios. Our main contributions can then be summarized as follows:
  • We propose a light-weight Test-Time Adaptation method for CLIP, called CLIPArTT, that adapts this VLM to new domains without a considerable computation overhead.
  • We introduce a new benchmark for Test-Time Adaptation on Vision-Language Models by implementing representative baselines, such as TENT.
  • Through comprehensive experiments, we subject our CLIPArTT methodology to diverse and challenging Test-Time Adaptation scenarios, each characterized by distinct types of domain shifts. The outcomes of these experiments highlight the performance of our approach when compared against other methodologies addressing similar challenges.
The rest of the paper is organized in the following sections. Section 2 compiles significant related work to contextualize our method. Section 3 introduces CLIPArTT and explains its functionality. Section 4 details the experimental setting, with results exposed in Section 5. Section 6 concludes with a summary of contributions and potential future work.
2 Related work
Test-Time Adaptation.  TTA is a particular setting of Domain Adaptation, encompassing two main characteristics: (a) adapting a model to a target domain, with inputs coming as unlabeled data streams (i.e., batches), without (b) any access to the source domain samples. The former challenge complicates accurately estimating the target domain’s distribution, while the latter impedes solving the problem by directly comparing domains’ characteristics such as statistics. Despite the challenging nature of the problem, the field has gained important momentum in recent years, providing insights on the possibilities and limitations of adapting pre-trained models.
An important branch of TTA methods targets batch normalization layers for adaptation, as they can keep key information from the source domain. As the simplest approach, PTBN Nado et al. (2021) addresses the problem by simply recomputing the batch statistics at test-time. TENT Wang et al. (2021) goes further by using entropy minimization on the model’s predictions as feedback, and optimizes the affine parameters in batch normalization layers after one iteration. Entropy minimization has since been a recurrent alternative to update models to target domains without label supervision Goyal et al. (2022); Zhang et al. (2022); Lin et al. (2023); Lee et al. (2023) by increasing the model’s confidence. These approaches, however, rely on image augmentations or large batches to be effective. More recently, methods such as Text-Prompt Tuning (TPT) Shu et al. (2022) have introduced entropy minimization to adapt CLIP, while focusing on learning an adapter for the text prompts. However, this method heavily relies on input augmentations, making it more costly to train. Our method also fine-tunes normalization layers to have a minimal impact on the model’s knowledge and utilizes the confidence of the predictions to improve the text supervision. Moreover, we do not employ any augmentations on the inputs.
Pseudo-labels have also played an important role in previous TTA research. SHOT Liang et al. (2020) employs them with cross-entropy to regularize mutual information between inputs and predictions. The entire feature encoder is optimized in this case. CoTTA Wang et al. (2022) makes use of pseudo-labels in the context of a student-teacher model, where one model processes the original input and the other processes augmented versions of that input. A consistency loss is used to compare both models’ outputs. PAD Wu et al. (2021) improves the quality of pseudo-labels by also augmenting inputs and using voting on the models’ predictions. Instead of simple class pseudo-labels, CLIPArTT improves the text supervision to broaden the likelihood of accurately predicting the right class.
Some other methods utilize less conventional approaches for TTA. LAME Boudiaf et al. (2022) refines the classifier’s predictions in a transductive manner using the similarity between features through the Laplacian. A family of test-time training (TTT) methods have also gained popularity Sun et al. (2020); Liu et al. (2021); Osowiechi et al. (2023); Gandelsman et al. (2022); Vargas Hakim et al. (2023); Osowiechi et al. (2024), where a sub-branch is trained alongside the main network in an unsupervised way, which is later used to update the model. These methods require training the model from scratch on the source domain. As LAME, our method is related to the concept of Laplacian regularization, however we exploit this concept very differently. Whereas LAME employs it to refine predictions in a traditional transductive inference approach where the model is kept as is, our method exploits in a test-time adaptation loss enforcing the consistency between embeddings of related batch samples. Contrary to TTT methods, we do not train the model from scratch with any additional branch.
Conformal Learning.  CLIPArTT is also related to the field of conformal learning, where intervals of confidence for new predictions are derived from previous experience Shafer and Vovk (2008). In a conformal prediction, a given level of certainty is assigned to a set $\mathcal{C}=\{c_{1},...,c_{K}\}$ with the $K$ most plausible classes that an input can belong to Angelopoulos and Bates (2021). We draw inspiration from this concept and build a conformal set of class predictions that can help adapting CLIP towards an accurate top--1 prediction. Our technique, however, stands out by not relying on image transformations such as in Shu et al. (2022); Lai et al. (2023) to filter out predictions.
3 Methodology
We start by presenting the vanilla CLIP model for classification and explain how it can be extended to test-time adaptation using entropy minimization. Building on the limitations of this approach, we then introduce our CLIPArTT method that leverages class uncertainty and the relationship between samples in a batch.
Figure 1: CLIPArTT pipeline overview: 1) Computing predictions from Image-Text Similarity, 2) generating a new text prompt by filtering the top-$K$ class predictions, 3) employing the new prompt to obtain image-to-image and text-to-text similarity scores, used to build a pseudo-label for cross-entropy.
3.1 CLIP-based classification
Contrastive Language-Image Pre-training (CLIP) Radford et al. (2021) consists of a visual encoder $f^{v}_{\theta}(\cdot)$, mapping an image $\mathbf{x}$ to visual features $\mathbf{z}^{v}\in{}^{D}$, and a text encoder $f^{t}_{\theta}(\cdot)$ transforming text prompts $\mathbf{t}$ to text features $\mathbf{z}^{t}\in{}^{D}$. The visual and text encoders are trained jointly with a contrastive loss so that the feature embeddings of training images and their corresponding text prompt are close to each other, while those of different training examples are pushed apart.
In a classification task with $K$ fixed classes, CLIP can be used to perform inference by encoding a pre-defined text prompt for each class, for example $\mathbf{t}_{k}=$a photo of a {class $k$}”. For a new image $\mathbf{x}_{i}$, the probability of belonging to class $k$ is then estimated based on cosine similarity,
$p_{ik}\,=\,\frac{\exp\big{(}\cos(\mathbf{z}^{v}_{i},\mathbf{z}^{t}_{k})/\tau\big{)}}{\sum_{j}\exp\big{(}\cos(\mathbf{z}^{v}_{i},\mathbf{z}^{t}_{j})/\tau\big{)}},\quad\cos(\mathbf{z},\mathbf{z}^{\prime})=\frac{{\mathbf{z}^{\top}}{\mathbf{z}^{\prime}}}{\|\mathbf{z}\|_{2}\!\cdot\!\|\mathbf{z}^{\prime}\|_{2}},$ (1)
where $\tau$ is a suitable softmax temperature.
The model in Eq. (1) can be used for test-time adaptation in various ways, the simplest one being entropy minimization as in TENT. This approach, which relies on the principle that the decision boundary lies in a low-density region of space giving rise to a low prediction entropy, adapts the model parameters by minimizing entropy on a test batch of size $B$:
$\mathcal{L}_{\mathrm{T\textsc{ent}}}(\theta)=-\frac{1}{B}\sum_{i=1}^{B}\sum_{k=1}^{K}p_{ik}\log p_{ik}$ (2)
However, this approach as well as similar ones based on pseudo-labels Liang et al. (2020); Wang et al. (2022); Wu et al. (2021) suffer from two important limitations. First, due to domain shifts, the model’s prediction may be unreliable (e.g., giving the highest probability to the wrong class) and techniques such as entropy minimization or standard pseudo-label will only reinforce these errors during adaptation. Secondly, they assume that samples in a test batch are independent and do not directly leverage their semantic relationships.
CIFAR10 CIFAR100
Top-1 Top-3 Top-1 Top-3
Original 88.74 97.79 61.68 80.92
C-Average 59.22 82.43 29.43 46.61
Table 1: Accuracy (%) on CIFAR-10/100 and CIFAR-10/100-C datasets with Level 5 corruption for the top-1 or the top-3 predicted classes.
Figure 2: Evolution of the accuracy after several iterations of adaptation at test-time for TENT.
3.2 Our CLIPArTT method
The proposed CLIPArTT method, illustrated in Figure 1, addresses the above-mentioned limitations using two key insights. The first insight, inspired by conformal learning, is that the correct class is often among the ones with the highest probabilities, although the model’s most confident prediction may not be always correct. This claim is supported by the results in Table 13 (Original), showing that the correct class is within the top-3 predictions 97.79% of the times for CIFAR10 (versus 88.74% within the top-1), and 80.92% of the times for CIFAR100 (versus 61.68% within the top-1). A way to include information about multiple classes could, therefore, help adapting the model at test time. The second insight is that the similarity between the batch samples could be evaluated based on their visual and/or text embeddings. As we will show below, such similarities could be exploited in a strategy related to graph-Laplacian regularization.
[…]
Transductive TTA  Next, we design a test-time adaptation loss that accounts for semantic relationships between batch samples. Let $\mathbf{Z}^{v}\in{}^{B\times D}$ and $\hat{\mathbf{Z}}^{t}\in{}^{B\times D}$ denote the normalized visual and instance-specific text embeddings of the samples within the test batch, respectively. We compute an image-to-image similarity matrix, $\mathbf{S}^{v}=\mathbf{Z}^{v}{(\mathbf{Z}^{v})^{\top}}\in[-1,1]^{B\times B}$, and a text-to-text similarity matrix, $\mathbf{S}^{t}=\hat{\mathbf{Z}}^{t}{(\hat{\mathbf{Z}}^{t})^{\top}}\in[-1,1]^{B\times B}$. The former measures the affinity between each pair of samples within the batch in terms of their visual characteristics (shapes, textures, etc.). The latter captures common (or related) classes in the top-$k$ predictions of two samples, since it is computed using the instance-specific multi-class prompts.
We deploy these two pairwise similarity matrices to compute pseudo-labels as follows:
$\mathbf{Q}=\text{softmax}\big{(}(\mathbf{S}^{v}+\mathbf{S}^{t})/2\tau\big{)}\in[0,1]^{B\times B}$ (3)
where the softmax operation is applied column-wise and the temperature $\tau=0.01$ in all our experiments.
Let $\hat{\mathbf{P}}$ denotes the zero-shot prediction matrix using our instance-specific multi-class text prompts:
$\hat{\mathbf{P}}=\text{softmax}\big{(}\mathbf{Z}^{v}{(\hat{\mathbf{Z}}^{t})^{\top}}/\tau\big{)},\quad\hat{p}_{ij}=\frac{\exp\big{(}\cos(\mathbf{z}^{v}_{i},\hat{\mathbf{z}}^{t}_{j})/\tau\big{)}}{\sum_{k}\exp\big{(}\cos(\mathbf{z}^{v}_{i},\hat{\mathbf{z}}^{t}_{k})/\tau\big{)}}$ (4)
This matrix, along with the pairwise pseudo-labels we introduced in Eq. (3), yield our final TTA loss based on the cross-entropy:
$\mathcal{L}_{\mathrm{TTA}}(\theta)=-\frac{1}{B}\sum_{i=1}^{B}\sum_{j=1}^{B}q_{ij}\log\hat{p}_{ij}.$ (5)
Unlike recent approaches Shu et al. (2022), which adapt the CLIP model by learning a text prompt, we instead update the normalization-layer parameters of the visual encoder, which yields a light-weight, computationally efficient TTA method.
3.3 Link to graph-Laplacian regularization
An important element of our TTA method is that it exploits the similarities between pairs of samples within the batch, in terms of the visual features and text prompts representing the top-$k$ classes. In this section, we draw a connection between the proposed method and graph-Laplacian regularization, well-known in the context semi-supervised learning techniques Chapelle et al. (2006) and label propagation Iscen et al. (2019). In our case, the Laplacian regularizer is computed over a set of $B$ nodes with predictions $\mathbf{Z}\in{}^{B\times D}$ as
$\mathcal{L}_{\mathrm{reg}}(\mathbf{Z})=\mathrm{trace}\big{(}{\mathbf{Z}^{\top}}\mathbf{L}_{\mathrm{W}}\mathbf{Z}\big{)}\,=\,\sum_{i,j}w_{ij}\|\mathbf{z}_{i}-\mathbf{z}_{j}\|_{2}^{2},$ (6)
where $\mathbf{L}_{\mathrm{W}}\in{}^{B\times B}$ is the Laplacian matrix defined from edge weight matrix $\mathbf{W}$ as follows:
$[\mathbf{L}_{\mathrm{W}}]_{ij}=\left\{\begin{array}[]{ll}d_{ii}=\sum_{j}w_{ij},&\text{ if }i=j\\ -w_{ij}&\text{ else}.\end{array}\right.$ (7)
This connection is made in the following proposition.
Proposition 1.
The TTA loss in Eq. (5) can be expressed as a Laplacian regularization over a bipartite graph with one set of nodes for image embeddings and another for text embeddings.
Proof.
Expanding the softmax in Eq. (5), we get
$\mathcal{L}_{\mathrm{TTA}}$ $=-\frac{1}{B}\sum_{i=1}^{B}\sum_{j=1}^{B}q_{ij}\log\frac{\exp\big{(}{(\mathbf{z}^{v}_{i})^{\top}}\mathbf{z}^{t}_{j}/\tau\big{)}}{\sum_{k}\exp\big{(}{(\mathbf{z}^{v}_{i})^{\top}}\mathbf{z}^{t}_{k}/\tau\big{)}}$ (8)
$=\frac{1}{\tau B}\sum_{i}\sum_{j}q_{ij}\Big{(}-\!{(\mathbf{z}^{v}_{i})^{\top}}\mathbf{z}^{t}_{j}\,+\,\tau\underbrace{\log\sum_{k}\exp\big{(}{(\mathbf{z}^{v}_{i})^{\top}}\mathbf{z}^{t}_{k}/\tau\big{)}}_{\text{LogSumExp (LSE)}}\Big{)}$ (9)
Since the feature embeddings are normalized, we have $\|\mathbf{z}_{i}^{v}\|_{2}=\|\mathbf{z}_{j}^{t}\|_{2}=1$ and thus ${(\mathbf{z}^{v}_{i})^{\top}}\mathbf{z}^{t}_{j}=2-\|\mathbf{z}^{v}_{i}-\mathbf{z}^{t}_{j}\|_{2}^{2}$. Moreover, following the scaled LogSumExp (LSE) rule, we can bound the right-side term as follows:
$\max_{k}\,\{{(\mathbf{z}^{v}_{i})^{\top}}\mathbf{z}^{t}_{k}\}\,<\,\tau LSE\,\leq\,\max_{k}\,\{{(\mathbf{z}^{v}_{i})^{\top}}\mathbf{z}^{t}_{k}\}\,+\,\tau\log\!B$ (10)
For a small $\tau$ (we use a value of 0.01 in our method), the bounds tighten and we get that
$LSE\,\approx\,\tfrac{1}{\tau}\max_{k}\,\{{(\mathbf{z}^{v}_{i})^{\top}}\mathbf{z}^{t}_{k}\}\,=\,\tfrac{1}{\tau}{(\mathbf{z}^{v}_{i})^{\top}}\mathbf{z}^{t}_{i}$ (11)
The last equality comes from the hypothesis that, in a well trained model, the maximum similarity occurs between the image and embedding of a samples and its corresponding text embedding. Our TTA loss can then be expressed as
$\mathcal{L}_{\mathrm{TTA}}$ $\approx\frac{1}{\tau B}\Big{(}\sum_{i}\sum_{j}q_{ij}\|\mathbf{z}^{v}_{i}-\mathbf{z}^{t}_{j}\|_{2}^{2}\ +\,\ \sum_{i}{(\mathbf{z}^{v}_{i})^{\top}}\mathbf{z}^{t}_{i}\Big{)}\,+\,\text{const}$ (12)
$=\frac{1}{\tau B}\mathrm{trace}\big{(}{(\mathbf{Z}^{v})^{\top}}\!(\mathbf{L}_{\mathrm{Q}}+\mathbf{I})\,\mathbf{Z}^{t}\big{)}\ +\,\text{const}$ (13)
where $\mathbf{I}$ is the identity matrix. The modified Laplacian $\mathbf{L}_{\mathrm{Q}}+\mathbf{I}$ enforces nodes with a high connection weight ($q_{ij}$) to have similar embeddings, while also avoiding embeddings to collapse into a single vector.
4 Experimental Settings
We evaluate CLIPArTT’s performance across diverse TTA datasets using established methodologies. These datasets simulate challenging domain shifts, offering insights into our approach’s efficacy. Additionally, we explore CLIPArTT’s adaptability on the original dataset through zero-shot test-time adaptation, crucial given the novel nature of the datasets. Our evaluation framework encompasses distinct categories, namely natural images, common corruptions, simulated shift and video shift, providing a comprehensive assessment of the model’s performance across diverse challenges.
In our evaluation of natural images, we utilize CIFAR-10, CIFAR-10.1, and CIFAR-100 datasets, each comprising 10,000 images and featuring 10 and 100 classes, respectively. These datasets represent natural imagery and are novel to the model under scrutiny. Notably, CIFAR-10.1 introduces a natural domain shift from CIFAR-10, thereby enriching our assessment with varied and nuanced data distributions.
Transitioning to our investigation of common corruptions, we turn to the CIFAR-10-C and CIFAR-100-C datasets Hendrycks and Dietterich (2019). These datasets offer a diverse range of 15 distinct corruptions, including elastic transform and impulse noise, among others. Each corruption is characterized by 5 severity levels, yielding a total of 75 unique testing scenarios per dataset. Within each severity level, there are 10,000 images, contributing to a comprehensive evaluation of the model’s robustness against a variety of corruption types and intensities.
Furthermore, we examine the challenging VisDA-C dataset Peng et al. (2018), which represents two domain shifts: the simulated shift and video shift. The former comprises a set of 152,397 images rendered in 3D across 12 different classes. The later includes 72,372 YouTube video frames spanning the same categories. This dataset presents an important challenge, as it bridges the gap of the type of imagery that a model can be applied on.
Test-time adaptation.  For test-time adaptation, model updates are applied to all the Layer Normalization (LN) layers within the visual encoder. We employ the ADAM optimizer with a fixed learning rate set to $10^{-3}$. Throughout our experiments, a consistent batch size of 128 is utilized to maintain uniformity and enable effective comparisons across different scenarios. As in previous TTA works, a smaller learning rate of $10^{-4}$ is preferred to adapt on the 3D renderings split Vargas Hakim et al. (2023), as it represents a more aggressive shift.
Benchmarking.  We compare the performance of CLIPArTT with state-of-the-art methods. Specifically, we utilize adapted versions of TENT Wang et al. (2021) and LAME Boudiaf et al. (2022) tailored to CLIP. The adapted TENT now updates only the affine parameters in Layer Normalization layers through entropy minimization directly on the similarity between image and text features, diverging from its original implementation (i.e., using classifier’s logits). Notably, we employ 10 iterations for TENT adaptation, a choice informed by our findings, as evidenced in Fig 2. For VisDA-C, we also explore using only 1 iteration, as in the original paper. Similarly, we utilize the image-to-text similarity scores as logits for LAME, where the Laplacian regularizer is applied to the similarity between image features. Finally, we extend the evaluation to include TPT Shu et al. (2022), a recent adaptation method for CLIP. TPT heavily relies on image augmentations, thus we reduce batch-size of 32 due to its high memory requirement.
$K$ = 1 $K$ = 3 $K$ = 4
Original 89.80 $\pm$0.05 90.04 $\pm$0.13 90.41 $\pm$0.07
CIFAR 10.1 85.37 $\pm$0.17 86.35 $\pm$0.27 86.07 $\pm$0.21
C-Average 70.79 71.17 70.99
Table 2: Accuracy (%) on CIFAR-10, CIFAR-10.1 and CIFAR-10-C datasets with Level 5 corruption for different number of $K$ selected classes to create pseudo-labels.
$K$ = 3 $K$ = 4 $K$ = 5
3D 84.31 84.24 87.24
YT 82.80 83.17 83.46
Table 3: Accuracy (%) on the VisDA-C 3D training split (3D) and validation split (YT) with ViT-B/32 on different numbers of selected classes ($K$) to create pseudo-labels.
5 Results
In this section, we outline the experimental outcomes derived from CLIPArTT and provide a comprehensive analysis. We first show a series of exploratory ablations that later help conduct the final experiments on the different datasets and compare with state-of-the-art.
5.1 Ablation Studies
Selection of the number of classes for new prompts.  Determining the number of classes in building new prompts presents a critical aspect in our methodology. Table 13 illustrates various scenarios where $K\!=\!3$ emerges as a favorable choice, exhibiting a remarkable accuracy of $100\%$ across multiple instances of CIFAR-10 datasets. This initial observation is further corroborated by the findings presented in Table 14. However, the decision becomes more nuanced when applied to CIFAR-100 datasets. While leveraging the top 3 classes contributes to enhanced accuracy, it does not guarantee optimal performance. Similarly, adopting a strategy akin to CIFAR-10, wherein $30\%$ of the classes are chosen, proves impractical due to resulting lengthy sentences that are impossible to tokenize. Examination of Table 4 reveals that the optimal choice varies depending on the distribution shift. Nonetheless, our analysis shows that $K\!=\!3$ consistently yields superior performance, particularly evident in the average results across CIFAR-100-C. Consequently, we maintain $K\!=\!3$ as the preferred configuration for all subsequent experiments on CIFAR datasets. A similar study was conducted for VisDA-C, in which $K\!=\!5$ (See Table 3) is chosen as the number of classes.
$K$ = 1 $K$ = 3 $K$ = 5 $K$ = 7 $K$ = 10 $K$ = 20
Original 69.00 $\pm$0.22 69.79 $\pm$0.04 69.68 $\pm$0.07 69.56 $\pm$0.02 69.78 $\pm$0.02 69.93 $\pm$0.08
Gaussian Noise 26.05 $\pm$0.11 25.32 $\pm$0.14 24.69 $\pm$0.03 24.60 $\pm$0.03 24.70 $\pm$0.15 23.73 $\pm$0.07
Shot noise 28.88 $\pm$0.11 27.90 $\pm$0.05 27.25 $\pm$0.26 26.75 $\pm$0.07 26.83 $\pm$0.26 25.73 $\pm$0.13
Impulse Noise 24.04 $\pm$0.09 25.62 $\pm$0.09 25.12 $\pm$0.14 25.25 $\pm$0.14 24.95 $\pm$0.23 24.57 $\pm$0.12
Defocus blur 49.03 $\pm$0.15 49.88 $\pm$0.23 49.75 $\pm$0.11 49.74 $\pm$0.25 49.62 $\pm$0.10 49.49 $\pm$0.07
Glass blur 26.77 $\pm$0.14 27.89 $\pm$0.03 27.76 $\pm$0.23 27.28 $\pm$0.07 26.57 $\pm$0.07 25.52 $\pm$0.09
Motion blur 46.50 $\pm$0.09 47.93 $\pm$0.14 47.48 $\pm$0.21 47.57 $\pm$0.10 47.53 $\pm$0.09 47.36 $\pm$0.09
Zoom blur 52.08 $\pm$0.12 52.70 $\pm$0.06 52.22 $\pm$0.10 52.10 $\pm$0.24 52.62 $\pm$0.24 52.82 $\pm$0.09
Snow 49.24 $\pm$0.07 49.72 $\pm$0.01 48.98 $\pm$0.08 48.87 $\pm$0.08 49.13 $\pm$0.08 49.54 $\pm$0.13
Frost 49.91 $\pm$0.07 49.63 $\pm$0.12 48.43 $\pm$0.17 48.11 $\pm$0.06 48.72 $\pm$0.08 49.13 $\pm$0.10
Fog 47.15 $\pm$0.04 48.77 $\pm$0.04 48.95 $\pm$0.18 48.78 $\pm$0.22 49.06 $\pm$0.05 48.74 $\pm$0.36
Brightness 60.01 $\pm$0.08 61.27 $\pm$0.08 60.77 $\pm$0.16 60.89 $\pm$0.19 60.98 $\pm$0.18 61.03 $\pm$0.19
Contrast 46.90 $\pm$0.21 48.55 $\pm$0.24 49.01 $\pm$0.14 49.07 $\pm$0.03 49.27 $\pm$0.09 49.08 $\pm$0.12
Elastic transform 36.32 $\pm$0.10 37.45 $\pm$0.08 37.63 $\pm$0.12 37.31 $\pm$0.09 37.13 $\pm$0.16 36.94 $\pm$0.11
Pixelate 32.52 $\pm$0.17 33.88 $\pm$0.14 34.40 $\pm$0.15 34.38 $\pm$0.16 34.65 $\pm$0.09 34.32 $\pm$0.02
JPEG compression 35.81 $\pm$0.11 36.07 $\pm$0.32 35.77 $\pm$0.01 35.60 $\pm$0.10 35.63 $\pm$0.10 35.29 $\pm$0.14
Average 40.75 41.51 41.21 41.09 41.16 40.89
Table 4: Accuracy (%) on CIFAR-100 and CIFAR-100-C datasets with Level 5 corruption for different number of $K$ selected classes to create pseudo-labels.
Selection of the number of iterations.  In line with our approach for TENT, we investigate optimal number of iterations. As demonstrated in Table 5, iterating 10 times strikes the most favorable balance, exhibiting superior performance across the overall average on CIFAR-10-C and yielding optimal results for numerous corruption types. Nonetheless, it is imperative to acknowledge that employing a lower number of iterations (e.g., Iter = 5) may yield improved outcomes in scenarios with minimal or no distribution shift, while a higher number of iterations (e.g., Iter = 20) may be more effective in mitigating severe distribution shifts. As a result, 10 iterations are used for all forthcoming experiments.
Iter = 1 Iter = 5 Iter = 10 Iter = 20
Original 89.59 $\pm$0.01 90.54 $\pm$0.09 90.04 $\pm$0.13 88.32 $\pm$0.12
CIFAR 10.1 84.78 $\pm$0.02 86.67 $\pm$0.06 86.35 $\pm$0.27 84.33 $\pm$0.31
Gaussian Noise 39.75 $\pm$0.04 53.79 $\pm$0.08 59.90 $\pm$0.36 59.33 $\pm$0.20
Shot noise 43.80 $\pm$0.04 57.16 $\pm$0.24 62.77 $\pm$0.07 63.08 $\pm$0.43
Impulse Noise 45.19 $\pm$0.07 52.44 $\pm$0.14 56.02 $\pm$0.16 56.73 $\pm$0.01
Defocus blur 72.93 $\pm$0.07 76.36 $\pm$0.10 76.74 $\pm$0.05 75.33 $\pm$0.13
Glass blur 46.61 $\pm$0.06 57.45 $\pm$0.13 61.77 $\pm$0.16 62.01 $\pm$0.27
Motion blur 67.89 $\pm$0.03 74.34 $\pm$0.14 76.01 $\pm$0.19 75.94 $\pm$0.28
Zoom blur 73.24 $\pm$0.05 77.03 $\pm$0.07 77.40 $\pm$0.20 75.42 $\pm$0.13
Snow 73.81 $\pm$0.07 76.51 $\pm$0.08 77.29 $\pm$0.16 76.18 $\pm$0.21
Frost 74.80 $\pm$0.04 78.13 $\pm$0.12 79.20 $\pm$0.08 77.44 $\pm$0.30
Fog 69.81 $\pm$0.06 74.16 $\pm$0.09 75.74 $\pm$0.14 74.66 $\pm$0.02
Brightness 84.16 $\pm$0.01 86.61 $\pm$0.10 86.59 $\pm$0.16 85.14 $\pm$0.42
Contrast 67.75 $\pm$0.03 74.85 $\pm$0.04 77.82 $\pm$0.14 77.75 $\pm$0.11
Elastic transform 63.15 $\pm$0.08 68.53 $\pm$0.19 70.20 $\pm$0.01 68.48 $\pm$0.24
Pixelate 54.20 $\pm$0.02 61.87 $\pm$0.04 66.52 $\pm$0.13 67.13 $\pm$0.15
JPEG compression 57.46 $\pm$0.09 62.00 $\pm$0.13 63.51 $\pm$0.14 63.64 $\pm$0.20
Average 62.30 68.75 71.17 70.55
Table 5: Accuracy (%) on CIFAR-10, CIFAR-10.1 and CIFAR-10-C datasets with Level 5 corruption for different number of iterations to update the model at test-time.
Pseudo-label selection for adaptation.  For the target, we employed a linear combination of both image-to-image and text-to-text similarities to encapsulate the information derived from both modalities. This combination proved to be more effective than each modality separately (see Table 6). While selecting solely the text-to-text similarity for adaptation performs well, it falls short of achieving optimal results. Conversely, utilizing only the image-to-image similarity mainly improves for the original CIFAR-10 dataset and few corruption types.
Image Text Image + Text
Original 90.18 $\pm$0.02 89.05 $\pm$0.14 90.04 $\pm$0.13
CIFAR 10.1 86.25 $\pm$0.37 84.85 $\pm$0.40 86.35 $\pm$0.27
Gaussian Noise 59.05 $\pm$0.30 59.29 $\pm$0.27 59.90 $\pm$0.36
Shot noise 62.11 $\pm$0.18 61.89 $\pm$0.31 62.77 $\pm$0.07
Impulse Noise 55.43 $\pm$0.12 55.48 $\pm$0.10 56.02 $\pm$0.16
Defocus blur 76.88 $\pm$0.09 76.25 $\pm$0.10 76.74 $\pm$0.05
Glass blur 61.56 $\pm$0.03 61.28 $\pm$0.33 61.77 $\pm$0.16
Motion blur 76.32 $\pm$0.15 75.37 $\pm$0.27 76.01 $\pm$0.19
Zoom blur 77.66 $\pm$0.09 76.29 $\pm$0.11 77.40 $\pm$0.20
Snow 77.28 $\pm$0.08 76.03 $\pm$0.16 77.29 $\pm$0.16
Frost 78.80 $\pm$0.18 78.05 $\pm$0.21 79.20 $\pm$0.08
Fog 75.69 $\pm$0.17 73.70 $\pm$0.15 75.74 $\pm$0.14
Brightness 86.62 $\pm$0.20 84.69 $\pm$0.04 86.59 $\pm$0.16
Contrast 77.43 $\pm$0.12 74.52 $\pm$0.02 77.82 $\pm$0.14
Elastic transform 69.63 $\pm$0.02 69.33 $\pm$0.20 70.20 $\pm$0.01
Pixelate 66.33 $\pm$0.16 64.86 $\pm$0.29 66.52 $\pm$0.13
JPEG compression 63.92 $\pm$0.13 63.44 $\pm$0.20 63.51 $\pm$0.14
Average 70.98 70.03 71.17
Table 6: Accuracy (%) on CIFAR-10, CIFAR-10.1 and CIFAR-10-C datasets with different targets.
CLIP BS = 8 BS = 16 BS = 32 BS = 64 BS = 128
Original 88.74 82.50 $\pm$0.13 85.89 $\pm$0.19 88.25 $\pm$0.15 89.48 $\pm$0.15 90.04 $\pm$0.13
CIFAR 10.1 83.25 77.2 $\pm$0.92 81.55 $\pm$0.53 84.00 $\pm$0.31 85.40 $\pm$0.08 86.35 $\pm$0.27
C-Average 59.22 61.10 64.72 67.70 69.82 71.17
Table 7: Accuracy (%) on CIFAR-10, CIFAR-10.1 and CIFAR-10-C datasets with Level 5 corruption for different number of batch-size.
Performance of the model for different batch-sizes.  TTA methods have historically exhibited limitations when applied with small batch sizes. In our study, we investigate the performance of our model in this regard. As depicted in Table 15, performance improves significantly with increasing batch size. However, beyond a batch size of 8, no further performance gains are observed. This phenomenon can be attributed to the fact that, with smaller batch sizes, our method has the potential to introduce uncertainty by introducing multiple classes, thereby diminishing performance, especially when the model is already confident. In subsequent experiments, we maintain a batch size of 128, consistent with prevailing practices in the state-of-the-art methodologies.
5.2 Experiments
CLIP LAME TENT TPT (BS=32) CLIPArTT
Original 88.74 89.36 $\pm$0.06 91.69 $\pm$0.10 88.06 $\pm$0.06 90.04 $\pm$0.13
CIFAR 10.1 83.25 81.22 $\pm$0.33 87.60 $\pm$0.45 81.80 $\pm$0.27 86.35 $\pm$0.27
Gaussian Noise 35.27 24.71 $\pm$0.11 41.27 $\pm$0.27 33.90 $\pm$0.08 59.90 $\pm$0.36
Shot noise 39.67 27.44 $\pm$0.09 47.20 $\pm$0.23 38.20 $\pm$0.02 62.77 $\pm$0.07
Impulse Noise 42.61 31.38 $\pm$0.15 48.58 $\pm$0.31 37.66 $\pm$0.20 56.02 $\pm$0.16
Defocus blur 69.76 62.45 $\pm$0.44 77.12 $\pm$0.16 67.83 $\pm$0.28 76.74 $\pm$0.05
Glass blur 42.40 29.96 $\pm$0.06 52.65 $\pm$0.30 38.81 $\pm$0.12 61.77 $\pm$0.16
Motion blur 63.97 54.00 $\pm$0.36 71.25 $\pm$0.09 63.39 $\pm$0.13 76.01 $\pm$0.19
Zoom blur 69.83 61.97 $\pm$0.36 76.20 $\pm$0.19 68.95 $\pm$0.16 77.40 $\pm$0.20
Snow 71.78 64.61 $\pm$0.48 78.29 $\pm$0.20 70.16 $\pm$0.10 77.29 $\pm$0.16
Frost 72.86 65.17 $\pm$0.17 79.84 $\pm$0.09 72.39 $\pm$0.22 79.20 $\pm$0.08
Fog 67.04 59.13 $\pm$0.49 77.39 $\pm$0.01 64.31 $\pm$0.28 75.74 $\pm$0.14
Brightness 81.87 80.05 $\pm$0.23 87.78 $\pm$0.03 81.30 $\pm$0.18 86.59 $\pm$0.16
Contrast 64.37 56.91 $\pm$0.37 79.47 $\pm$0.11 62.26 $\pm$0.31 77.82 $\pm$0.14
Elastic transform 60.83 53.89 $\pm$0.20 70.00 $\pm$0.25 56.43 $\pm$0.27 70.20 $\pm$0.01
Pixelate 50.53 39.67 $\pm$0.34 63.74 $\pm$0.18 42.80 $\pm$0.40 66.52 $\pm$0.13
JPEG compression 55.48 47.24 $\pm$0.14 62.64 $\pm$0.14 53.67 $\pm$0.25 63.51 $\pm$0.14
Average 59.22 50.57 67.56 56.80 71.17
Table 8: Accuracy (%) on CIFAR-10, CIFAR-10.1 and CIFAR-10-C datasets with ViT-B/32 as visual encoder.
Performance analysis under natural or no domain shift.  In the different Tables 8, 9, 10, and 11, CLIPArTT consistently enhances accuracy across CIFAR-10, CIFAR10.1, and CIFAR-100 datasets compared to the baseline (+2%, +3%, +8% respectively with ViT-B/32). This improvement suggests that even datasets devoid of corruption, which may not have been seen by the model during training, can benefit from adaptation. Consequently, a zero-shot model can leverage adaptation at test-time, relying solely on the model’s predictions. However, while CLIPArTT outperforms the baseline, as well as LAME and TPT with this fixed version, it is worth noting that TENT may yield superior results depending on the visual encoder and the number of classes. Entropy minimization used solely on confident results proves to be also effective. Interestingly, the corruptions in which TENT achieves a better performance tend to already have higher accuracy values.
CLIP LAME TENT TPT (BS=32) CLIPArTT
Original 61.68 58.27 $\pm$0.17 69.74 $\pm$0.16 63.78 $\pm$0.28 69.79 $\pm$0.04
Gaussian Noise 14.80 12.72 $\pm$0.04 14.38 $\pm$0.14 14.03 $\pm$0.10 25.32 $\pm$0.14
Shot noise 16.03 13.78 $\pm$0.08 17.34 $\pm$0.27 15.25 $\pm$0.17 27.90 $\pm$0.05
Impulse Noise 13.85 07.82 $\pm$0.14 10.03 $\pm$0.13 13.01 $\pm$0.13 25.62 $\pm$0.09
Defocus blur 36.74 33.38 $\pm$0.11 49.05 $\pm$0.07 37.60 $\pm$0.17 49.88 $\pm$0.23
Glass blur 14.19 09.00 $\pm$0.05 03.71 $\pm$0.07 16.41 $\pm$0.02 27.89 $\pm$0.03
Motion blur 36.14 32.79 $\pm$0.13 46.62 $\pm$0.27 37.52 $\pm$0.23 47.93 $\pm$0.14
Zoom blur 40.24 37.57 $\pm$0.15 51.84 $\pm$0.15 42.99 $\pm$0.11 52.70 $\pm$0.06
Snow 38.95 35.49 $\pm$0.18 46.71 $\pm$0.21 42.35 $\pm$0.13 49.72 $\pm$0.01
Frost 40.56 37.22 $\pm$0.21 44.90 $\pm$0.27 43.31 $\pm$0.14 49.63 $\pm$0.12
Fog 38.00 35.94 $\pm$0.09 47.31 $\pm$0.04 38.81 $\pm$0.17 48.77 $\pm$0.04
Brightness 48.18 44.93 $\pm$0.08 60.58 $\pm$0.18 50.23 $\pm$0.11 61.27 $\pm$0.08
Contrast 29.53 27.52 $\pm$0.06 45.90 $\pm$0.11 28.09 $\pm$0.09 48.55 $\pm$0.24
Elastic transform 26.33 24.01 $\pm$0.02 33.09 $\pm$0.08 28.12 $\pm$0.15 37.45 $\pm$0.08
Pixelate 21.98 19.55 $\pm$0.13 26.47 $\pm$0.09 20.43 $\pm$0.14 33.88 $\pm$0.14
JPEG compression 25.91 21.77 $\pm$0.14 29.89 $\pm$0.07 28.82 $\pm$0.09 36.07 $\pm$0.32
Average 29.43 26.23 35.19 30.46 41.51
Table 9: Accuracy (%) on CIFAR-100 and CIFAR-100-C datasets with ViT-B/32 as visual encoder.
Performance analysis under common corruptions.  In the various referenced tables, including Tables 8, 9, 10, and 11, CLIPArTT consistently outperforms the baseline, TPT, and LAME across all visual encoders, resulting in enhancements of up to 13% on average. A more detailed examination reveals that CLIPArTT frequently surpasses TENT, particularly in instances where the baseline performs poorly. This observation implies that CLIPArTT consistently outperforms TENT, which tends to compromise the model’s performance under conditions of low baseline confidence. For instance, in Table 9, TENT experiences a 3% decrease compared to the baseline under Impulse Noise, while CLIPArTT exhibits a 12% improvement. Conversely, when the baseline is confident, CLIPArTT consistently maintains a high performance. As an illustration, Table 8 shows that while CLIP achieves an accuracy of 81.87% under Brightness corruption, both TENT and CLIPArTT contribute to improvements of 7% and 6%, respectively.
CLIP TENT TPT (BS=32) CLIPArTT
Original ViT-B/16 89.25 92.75 $\pm$0.17 89.80 $\pm$0.05 92.61 $\pm$0.05
ViT-L/14 95.36 96.13 $\pm$0.06 95.18 $\pm$0.02 95.16 $\pm$0.03
CIFAR-10.1 ViT-B/16 84.00 88.52 $\pm$0.33 83.75 $\pm$0.21 88.72 $\pm$0.33
ViT-L/14 91.20 92.22 $\pm$0.25 91.32 $\pm$0.12 91.02 $\pm$0.02
C-Average ViT-B/16 60.15 68.00 59.75 73.22
ViT-L/14 76.04 79.18 75.01 78.06
Table 10: Accuracy (%) on CIFAR-10, CIFAR-10.1 and CIFAR-10-C datasets with ViT-B/16 and ViT-L/14 as visual encoders.
CLIP TENT TPT (BS=32) CLIPArTT
Original ViT-B/16 64.76 71.73 $\pm$0.14 67.15 $\pm$0.23 71.34 $\pm$0.07
ViT-L/14 73.28 78.03 $\pm$0.08 76.85 $\pm$0.06 79.42 $\pm$0.08
C-Average ViT-B/16 32.01 37.90 33.73 40.08
ViT-L/14 44.59 50.14 47.58 52.52
Table 11: Accuracy (%) on CIFAR-100 and CIFAR-100-C datasets with ViT-B/16 and ViT-L/14 as visual encoders.
CLIP TENT TENT-10 LAME TPT (BS=32) CLIPArTT
3D ViT-B/16 87.04 87.15 $\pm$0.21 88.08 $\pm$0.05 85.83 $\pm$0.06 83.85 $\pm$0.06 88.01 $\pm$0.47
ViT-B/32 84.51 84.75 $\pm$0.17 85.82 $\pm$0.02 83.85 $\pm$0.03 78.45 $\pm$0.05 87.24 $\pm$0.04
ViT-L/14 90.62 90.93 $\pm$0.01 91.33 $\pm$0.01 90.15 $\pm$0.07 91.07 $\pm$0.05 91.18 $\pm$0.01
YT ViT-B/16 86.71 86.96 $\pm$0.02 87.44 $\pm$0.02 86.24 $\pm$0.0 85.69 $\pm$0.05 86.22 $\pm$0.01
ViT-B/32 84.45 84.72 $\pm$0.01 85.46 $\pm$0.02 84.41 $\pm$0.02 82.95 $\pm$0.01 83.46 $\pm$0.03
ViT-L/14 84.93 85.38 $\pm$0.01 86.65 $\pm$0.05 85.08 $\pm$0.04 85.54 $\pm$0.38 85.55 $\pm$0.01
Table 12: Accuracy (%) on the VisDA-C training split (3D) and validation split (YT) with the three different feature extractors.
Performance analysis under simulated and video shifts.  Results on the 3D and YT splits of VisDA-C are shown in Table 12. With respect to pure CLIP, our method achieves an important improvement of accuracy with the three different image encoders. CLIPArTT also achieves a competitive advantage against TENT, LAME and TPT. We observed however that using TENT with 10 iterations (here named TENT-10) is a strong contender, achieving practically the same score on 3D images and outperforming on YouTube frames by a small margin. This can mean that TENT’s entropy minimization can achieve faster improvement with certain domain shifts.
6 Conclusions
In this study, we introduce CLIPArTT, the first comprehensive Test-Time Adaptation framework designed specifically for Vision-Language Models. By leveraging the model’s predictions as new pseudo-labels, we effectively minimize cross-entropy loss at test-time, thereby enhancing model performance even within a zero-shot setting. To ensure fair comparisons, we also incorporate Test-Time Adaptation models into this new benchmark for VLMs.
A comprehensive ablation study is conducted to determine optimal hyperparameters and gain deeper insights into the various model configurations. Our experimental results demonstrate that CLIPArTT achieves highly competitive performance across TTA datasets, surpassing state-of-the-art approaches. While TENT remains a strong competitor, our model offers enhanced versatility by effectively addressing both natural and severe domain shifts, thus exhibiting robustness across various scenarios.
Exploring the potential of text prompts in classification represents a promising avenue for future research. Investigating alternative methods to fine-tune these text prompts could yield valuable insights. Moreover, extending the study of Test-Time Adaptation to other scenarios, such as segmentation or object detection with Vision-Language Models (VLMs), holds significant promise for advancing our understanding of model adaptability and performance across diverse tasks.
CLIPArTT: Light-weight Adaptation of CLIP to New Domains at Test Time – Supplementary Material
Appendix A Unsupervised clustering
In Fig. 3, tSNE visualizations of data points are shown. We show how the distribution of data points change after adaptation, which improves the accuracy of class predictions and facilitates the assignment of ground truth labels.
Refer to caption Refer to caption
(a) Prediction (before adaptation) (b) Prediction (after adaptation)
Refer to caption Refer to caption
(c) Ground truth (before adaptation) (d) Ground truth (after adaptation)
Figure 3: The t-SNE visualizations exhibit discernible attributes of brightness within the visual features derived from CLIPArTT. Panels (a) and (b) present the model’s predictions before and after 10 iterations of adaptation, respectively. Panels (c) and (d) demonstrate the actual labels in the absence of adaptation and following adaptation of the representations, respectively.
Appendix B Evolution of similarity matrices
A fundamental component of CLIPArTT is the employment of image-to-image and text-to-text similarity matrices to build new pseudo labels. In Fig. 4, we show an example of a batch of five images. While the largest image-to-image similarity values concentrate on the main diagonal, the text-to-text similarity using the $K$ most confident classes for the new text prompt, highly impacts the correlation between images and text (CLIPArTT logits).
Figure 4: Similarity matrices before (a) and after (b) adaptation using $K=5$ classes. There is an increase in the confidence of the logits when using a pseudolabel from the new prompt with the $K$ most confident classes.
Appendix C Comprehensive experimental results
We present comprehensive tables containing all the detailed information about results that was summarized in the main paper.
CIFAR10 CIFAR100
Top 1 Top 3 Top 1 Top 3
ORIGINAL 88.74 100.00 61.68 97.34
Gaussian Noise 35.27 99.87 14.8 63.66
Shot noise 39.67 99.99 16.03 67.02
Impulse Noise 42.61 100.00 13.85 64.4
Defocus blur 69.76 100.00 36.74 90.14
Glass blur 42.40 100.00 14.19 61.66
Motion blur 63.97 100.00 36.14 90.36
Zoom blur 69.83 100.00 40.24 91.27
Snow 71.78 100.00 38.95 91.40
Frost 72.86 100.00 40.56 92.23
Fog 67.04 99.98 38.00 91.51
Brightness 81.87 100.00 48.18 93.10
Contrast 64.37 100.00 29.53 84.67
Elastic transform 60.83 100.00 26.33 78.96
Pixelate 50.53 100.00 21.98 75.65
JPEG compression 55.48 100.00 25.91 80.81
Average 59.22 99.99 29.43 81.12
Table 13: Accuracy (%) on CIFAR-10/100 and CIFAR-10/100-C datasets with Level 5 corruption for the top 1 or the top 3 predicted classes.
K = 1 K = 3 K = 4
ORIGINAL 89.8 $\pm$0.05 90.04 $\pm$0.13 90.41 $\pm$0.07
CIFAR 10.1 85.37 $\pm$0.17 86.35 $\pm$0.27 86.07 $\pm$0.21
Gaussian Noise 60.2 $\pm$0.24 59.90 $\pm$0.36 59.71 $\pm$0.15
Shot noise 62.08 $\pm$0.11 62.77 $\pm$0.07 62.17 $\pm$0.16
Impulse Noise 54.33 $\pm$0.07 56.02 $\pm$0.16 56.27 $\pm$0.15
Defocus blur 77.16 $\pm$0.02 76.74 $\pm$0.05 76.79 $\pm$0.11
Glass blur 61.91 $\pm$0.15 61.77 $\pm$0.16 61.72 $\pm$0.23
Motion blur 74.94 $\pm$0.15 76.01 $\pm$0.19 76.33 $\pm$0.1
Zoom blur 76.84 $\pm$0.13 77.40 $\pm$0.20 77.15 $\pm$0.04
Snow 76.87 $\pm$0.05 77.29 $\pm$0.16 76.56 $\pm$0.16
Frost 77.81 $\pm$0.04 79.20 $\pm$0.08 78.42 $\pm$0.04
Fog 75.83 $\pm$0.28 75.74 $\pm$0.14 75.65 $\pm$0.06
Brightness 85.55 $\pm$0.12 86.59 $\pm$0.16 86.83 $\pm$0.1
Contrast 78.02 $\pm$0.18 77.82 $\pm$0.14 78.27 $\pm$0.14
Elastic transform 69.42 $\pm$0.07 70.20 $\pm$0.01 69.81 $\pm$0.2
Pixelate 66.07 $\pm$0.09 66.52 $\pm$0.13 66.45 $\pm$0.08
JPEG compression 64.82 $\pm$0.26 63.51 $\pm$0.14 62.72 $\pm$0.25
Average 70.79 71.17 70.99
Table 14: Accuracy (%) on CIFAR-10, CIFAR-10.1 and CIFAR-10-C datasets with Level 5 corruption for different number of K selected classes to create pseudo-label.
CLIP BS = 8 BS = 16 BS = 32 BS = 64 BS = 128
ORIGINAL 88.74 82.50 $\pm$0.13 85.89 $\pm$0.19 88.25 $\pm$0.15 89.48 $\pm$0.15 90.04 $\pm$0.13
CIFAR 10.1 83.25 77.2 $\pm$0.92 81.55 $\pm$0.53 84.00 $\pm$0.31 85.40 $\pm$0.08 86.35 $\pm$0.27
Gaussian Noise 35.27 47.30 $\pm$0.37 50.91 $\pm$0.35 54.23 $\pm$0.28 57.89 $\pm$0.13 59.90 $\pm$0.36
Shot noise 39.67 49.62 $\pm$0.26 53.1 $\pm$0.27 56.88 $\pm$0.23 60.56 $\pm$0.12 62.77 $\pm$0.07
Impulse Noise 42.61 47.24 $\pm$0.22 50.24 $\pm$0.48 52.7 $\pm$0.21 54.88 $\pm$0.17 56.02 $\pm$0.16
Defocus blur 69.76 68.24 $\pm$0.35 72.22 $\pm$0.04 75.09 $\pm$0.16 75.97 $\pm$0.27 76.74 $\pm$0.05
Glass blur 42.40 49.49 $\pm$0.30 53.27 $\pm$0.04 57.18 $\pm$0.24 60.12 $\pm$0.14 61.77 $\pm$0.16
Motion blur 63.97 65.22 $\pm$0.06 69.02 $\pm$0.30 72.54 $\pm$0.27 74.71 $\pm$0.18 76.01 $\pm$0.19
Zoom blur 69.83 67.69 $\pm$0.20 71.33 $\pm$0.11 74.53 $\pm$0.11 76.35 $\pm$0.07 77.40 $\pm$0.20
Snow 71.78 68.68 $\pm$0.42 72.37 $\pm$0.11 74.93 $\pm$0.18 76.53 $\pm$0.41 77.29 $\pm$0.16
Frost 72.86 70.35 $\pm$0.25 73.93 $\pm$0.34 76.81 $\pm$0.23 78.22 $\pm$0.13 79.20 $\pm$0.08
Fog 67.04 66.25 $\pm$0.31 69.71 $\pm$0.24 72.36 $\pm$0.23 73.96 $\pm$0.21 75.74 $\pm$0.14
Brightness 81.87 77.36 $\pm$0.17 81.20 $\pm$0.20 84.07 $\pm$0.08 85.58 $\pm$0.25 86.59 $\pm$0.16
Contrast 64.37 65.12 $\pm$0.07 69.02 $\pm$0.12 72.60 $\pm$0.46 75.79 $\pm$0.24 77.82 $\pm$0.14
Elastic transform 60.83 59.61 $\pm$0.11 63.67 $\pm$0.13 66.36 $\pm$0.26 68.74 $\pm$0.07 70.20 $\pm$0.01
Pixelate 50.53 56.78 $\pm$0.24 60.01 $\pm$0.06 62.57 $\pm$0.19 64.64 $\pm$0.03 66.52 $\pm$0.13
JPEG compression 55.48 57.59 $\pm$0.26 60.78 $\pm$0.12 62.63 $\pm$0.06 63.43 $\pm$0.16 63.51 $\pm$0.14
Average 59.22 61.10 64.72 67.70 69.82 71.17
Table 15: Accuracy (%) on CIFAR-10, CIFAR-10.1 and CIFAR-10-C datasets with Level 5 corruption for different batch sizes.
CLIP TENT TPT (BS=32) CLIPArTT
original 89.25 92.75 $\pm$0.17 89.80 $\pm$0.05 92.61 $\pm$0.05
CIFAR 10.1 84.00 88.52 $\pm$0.33 83.75.0.21 $\pm$ 88.72 $\pm$0.33
Gaussian Noise 37.75 31.04 $\pm$0.38 35.35 $\pm$0.15 60.89 $\pm$0.26
Shot noise 41.10 40.54 $\pm$0.41 41.03 $\pm$0.19 65.19 $\pm$0.21
Impulse Noise 51.71 58.03 $\pm$0.16 54.86 $\pm$0.07 67.55 $\pm$0.09
Defocus blur 70.07 77.57 $\pm$0.03 70.29 $\pm$0.02 78.92 $\pm$0.12
Glass blur 42.24 47.16 $\pm$0.05 37.86 $\pm$0.17 57.18 $\pm$0.20
Motion blur 65.81 76.16 $\pm$0.05 67.43 $\pm$0.11 76.59 $\pm$0.06
Zoom blur 72.50 79.64 $\pm$0.12 72.91 $\pm$0.02 79.62 $\pm$0.11
Snow 73.23 81.68 $\pm$0.03 72.98 $\pm$0.32 81.13 $\pm$0.29
Frost 76.52 83.22 $\pm$0.05 75.87 $\pm$0.16 81.24 $\pm$0.08
Fog 68.35 80.78 $\pm$0.15 69.13 $\pm$0.27 78.47 $\pm$0.19
Brightness 83.36 89.85 $\pm$0.11 83.67 $\pm$0.14 88.66 $\pm$0.15
Contrast 61.90 79.24 $\pm$0.19 62.16 $\pm$0.06 75.15 $\pm$0.07
Elastic transform 53.16 62.54 $\pm$0.08 51.26 $\pm$0.23 69.49 $\pm$0.08
Pixelate 48.48 67.08 $\pm$0.24 44.65 $\pm$0.21 71.80 $\pm$0.16
JPEG compression 56.05 65.42 $\pm$0.05 56.73 $\pm$0.07 66.42 $\pm$0.25
Average 60.15 68.00 59.75 73.22
Table 16: Accuracy (%) on CIFAR-10, CIFAR-10.1 and CIFAR-10-C datasets with ViT-B/16 as visual encoder.
CLIP TENT TPT (BS=32) CLIPArTT
original 95.36 96.13 $\pm$0.06 95.18 $\pm$0.02 95.16 $\pm$0.03
CIFAR 10.1 91.20 92.22 $\pm$0.25 91.32 $\pm$0.12 91.02 $\pm$0.02
Gaussian Noise 64.64 68.87 $\pm$0.20 64.44 $\pm$0.11 70.04 $\pm$0.31
Shot noise 67.82 71.95 $\pm$0.06 66.81 $\pm$0.19 71.44 $\pm$0.16
Impulse Noise 78.21 80.22 $\pm$0.19 76.46 $\pm$0.17 79.42 $\pm$0.15
Defocus blur 80.73 83.10 $\pm$0.03 79.01 $\pm$0.23 81.75 $\pm$0.19
Glass blur 50.29 57.12 $\pm$0.07 49.64 $\pm$0.23 58.13 $\pm$0.23
Motion blur 80.75 82.69 $\pm$0.11 78.85 $\pm$0.04 80.76 $\pm$0.12
Zoom blur 82.75 84.91 $\pm$0.08 82.32 $\pm$0.13 83.39 $\pm$0.05
Snow 83.01 85.99 $\pm$0.11 82.69 $\pm$0.10 84.48 $\pm$0.07
Frost 84.90 87.15 $\pm$0.12 84.63 $\pm$0.08 85.21 $\pm$0.06
Fog 78.44 81.30 $\pm$0.07 77.56 $\pm$0.17 79.27 $\pm$0.07
Brightness 91.67 93.07 $\pm$0.04 90.94 $\pm$0.04 91.87 $\pm$0.09
Contrast 84.20 87.93 $\pm$0.04 82.88 $\pm$0.09 86.19 $\pm$0.06
Elastic transform 65.45 69.96 $\pm$0.12 64.81 $\pm$0.14 67.43 $\pm$0.24
Pixelate 75.10 77.89 $\pm$0.05 72.92 $\pm$0.12 77.11 $\pm$0.10
JPEG compression 72.58 75.49 $\pm$0.07 71.18 $\pm$0.19 74.46 $\pm$0.11
Average 76.04 79.18 75.01 78.06
Table 17: Accuracy (%) on CIFAR-10, CIFAR-10.1 and CIFAR-10-C datasets with ViT-L/14 as visual encoder.
CLIP TENT TPT (BS=32) CLIPArTT
original 64.76 71.73 $\pm$0.14 67.15 $\pm$0.23 71.34 $\pm$0.07
Gaussian Noise 15.88 12.28 $\pm$0.20 15.43 $\pm$0.03 19.01 $\pm$0.24
Shot noise 17.49 15.07 $\pm$0.21 16.88 $\pm$0.07 20.27 $\pm$0.21
Impulse Noise 21.43 13.13 $\pm$0.16 22.12 $\pm$0.15 17.66 $\pm$0.10
Defocus blur 40.10 50.35 $\pm$0.03 41.08 $\pm$0.22 49.86 $\pm$0.13
Glass blur 13.48 04.84 $\pm$0.14 18.43 $\pm$0.15 18.34 $\pm$0.31
Motion blur 39.82 49.85 $\pm$0.37 40.85 $\pm$0.26 50.00 $\pm$0.09
Zoom blur 45.45 54.76 $\pm$0.04 46.77 $\pm$0.06 54.13 $\pm$0.08
Snow 42.77 52.38 $\pm$0.18 47.24 $\pm$0.18 52.80 $\pm$0.27
Frost 45.39 51.66 $\pm$0.04 48.61 $\pm$0.14 49.56 $\pm$0.08
Fog 38.98 50.74 $\pm$0.14 39.92 $\pm$0.16 49.92 $\pm$0.11
Brightness 52.55 64.26 $\pm$0.09 55.83 $\pm$0.10 63.76 $\pm$0.13
Contrast 33.32 48.69 $\pm$0.08 33.13 $\pm$0.16 47.86 $\pm$0.02
Elastic transform 24.39 33.56 $\pm$0.28 27.36 $\pm$0.10 32.93 $\pm$0.23
Pixelate 21.89 36.20 $\pm$0.28 21.26 $\pm$0.10 39.49 $\pm$0.21
JPEG compression 27.21 30.80 $\pm$0.05 30.97 $\pm$0.10 35.56 $\pm$0.23
Average 32.01 37.90 33.73 40.08
Table 18: Accuracy (%) on CIFAR-100 and CIFAR-100-C datasets with ViT-B/16 as visual encoder.
CLIP TENT TPT (BS=16) CLIPArTT
ORIGINAL 73.28 78.03 $\pm$0.08 76.85 $\pm$0.06 79.42 $\pm$0.08
Gaussian Noise 30.55 36.93 $\pm$0.03 36.10 $\pm$0.11 41.46 $\pm$0.15
Shot noise 34.58 40.96 $\pm$0.16 38.23 $\pm$0.13 44.27 $\pm$0.09
Impulse Noise 44.89 49.09 $\pm$0.14 49.69 $\pm$0.21 51.44 $\pm$0.23
Defocus blur 48.88 55.23 $\pm$0.07 50.43 $\pm$0.19 56.55 $\pm$0.22
Glass blur 23.46 27.02 $\pm$0.23 24.35 $\pm$0.22 30.47 $\pm$0.14
Motion blur 50.83 56.03 $\pm$0.20 51.94 $\pm$0.04 56.98 $\pm$0.18
Zoom blur 56.02 61.19 $\pm$0.10 56.96 $\pm$0.16 62.56 $\pm$0.04
Snow 49.03 55.60 $\pm$0.09 54.89 $\pm$0.11 58.81 $\pm$0.11
Frost 53.27 58.21 $\pm$0.15 58.15 $\pm$0.33 60.38 $\pm$0.23
Fog 48.51 53.37 $\pm$0.25 49.26 $\pm$0.13 54.38 $\pm$0.04
Brightness 60.53 67.34 $\pm$0.17 66.60 $\pm$0.10 69.63 $\pm$0.14
Contrast 50.24 59.91 $\pm$0.13 53.64 $\pm$0.24 63.39 $\pm$0.13
Elastic transform 35.07 38.49 $\pm$0.12 35.72 $\pm$0.09 39.57 $\pm$0.39
Pixelate 43.86 48.37 $\pm$0.17 44.32 $\pm$0.10 50.45 $\pm$0.16
JPEG compression 39.11 44.42 $\pm$0.09 43.44 $\pm$0.11 47.45 $\pm$0.14
Average 44.59 50.14 47.58 52.52
Table 19: Accuracy (%) on CIFAR-100 and CIFAR-100-C datasets with ViT-L/14 as visual encoder.

### v2

@@
-CLIPArTT: Light-weight Adaptation of CLIP to New Domains at Test Time
+CLIPArTT: Adaptation of CLIP to New Domains at Test Time
 Abstract
@@
 Additionally, we pioneer the standardization of TTA benchmarks (e.g., TENT) in the realm of VLMs.
-Our findings demonstrate that, without requiring additional transformations nor new trainable modules, CLIPArTT enhances performance dynamically across non-corrupted datasets such as CIFAR-10, corrupted datasets like CIFAR-10-C and CIFAR-10.1, alongside synthetic datasets such as VisDA-C.
+Our findings demonstrate that, without requiring additional transformations nor new trainable modules, CLIPArTT enhances performance dynamically across non-corrupted datasets such as CIFAR-100, corrupted datasets like CIFAR-100-C and ImageNet-C, alongside synthetic datasets such as VisDA-C.
 This research underscores the potential for improving VLMs’ adaptability through novel test-time strategies, offering insights for robust performance across varied datasets and environments.
@@
 Additionally, we introduce a powerful adaptation technique that achieves state-of-the-art performance without a significant computational overhead.
-Comprehensive studies are performed on three datasets containing different types of domain shifts on several levels of severity, resulting in a total of 155 evaluation scenarios.
-Our main contributions can then be summarized as follows:
-  • We propose a light-weight Test-Time Adaptation method for CLIP, called CLIPArTT, that adapts this VLM to new domains without a considerable computation overhead.
+Comprehensive studies are performed on multiple datasets containing different types of domain shifts on several levels of severity, resulting in a total of numerous evaluation scenarios.
+Our main contributions can be summarized as follows:
+  • We propose a Test-Time Adaptation method for CLIP, named CLIPArTT, which adapts the VLM by updating normalization-layer parameters.
+This is achieved by combining multiple classes into a single new text prompt, which is then used as a pseudo-label.
   • We introduce a new benchmark for Test-Time Adaptation on Vision-Language Models by implementing representative baselines, such as TENT.
@@
 The outcomes of these experiments highlight the performance of our approach when compared against other methodologies addressing similar challenges.
-The rest of the paper is organized in the following sections.
-Section 2 compiles significant related work to contextualize our method.
-Section 3 introduces CLIPArTT and explains its functionality.
-Section 4 details the experimental setting, with results exposed in Section 5.
-Section 6 concludes with a summary of contributions and potential future work.
 2 Related work
@@
 Despite the challenging nature of the problem, the field has gained important momentum in recent years, providing insights on the possibilities and limitations of adapting pre-trained models.
-An important branch of TTA methods targets batch normalization layers for adaptation, as they can keep key information from the source domain.
-As the simplest approach, PTBN Nado et al.
-(2021) addresses the problem by simply recomputing the batch statistics at test-time.
-TENT Wang et al.
-(2021) goes further by using entropy minimization on the model’s predictions as feedback, and optimizes the affine parameters in batch normalization layers after one iteration.
-Entropy minimization has since been a recurrent alternative to update models to target domains without label supervision Goyal et al.
+A key focus in TTA methods is adapting batch normalization layers, which retain important source domain information.
+PTBN Nado et al.
+(2021) adjusts batch statistics at test time, while TENT Wang et al.
+(2021) refines affine parameters using entropy minimization on predictions.
+Entropy minimization, used in various methods Goyal et al.
 (2022); Zhang et al.
@@
 (2023); Lee et al.
-(2023) by increasing the model’s confidence.
-These approaches, however, rely on image augmentations or large batches to be effective.
-More recently, methods such as Text-Prompt Tuning (TPT) Shu et al.
-(2022) have introduced entropy minimization to adapt CLIP, while focusing on learning an adapter for the text prompts.
-However, this method heavily relies on input augmentations, making it more costly to train.
-Our method also fine-tunes normalization layers to have a minimal impact on the model’s knowledge and utilizes the confidence of the predictions to improve the text supervision.
-Moreover, we do not employ any augmentations on the inputs.
-Pseudo-labels have also played an important role in previous TTA research.
+(2023), enhances model confidence without label supervision but often depends on image augmentations or large batches.
+Text-Prompt Tuning (TPT) Shu et al.
+(2022) applies entropy minimization to CLIP by learning text prompt adapters, though it is costly due to required augmentations.
+Test-Time Distribution Normalization (TTDN) Zhou et al.
+(2023) normalizes test data to match the training distribution but needs access to source data or approximations.
+Our method fine-tunes normalization layers, leverages prediction confidence for text supervision, and avoids input augmentations.
+Recent techniques have also sought to efficiently adapt CLIP in a gradient-free manner.
+CALIP Guo et al.
+(2023) adapts the visual and text features bidirectionally through a parametric-free attention module, and later combine them to obtain new logits.
+Although efficient, this method requires the features of the entire test set to perform a hyperparameter search, which limits its generalization to very large datasets.
+TDA Karmanov et al.
+(2024) dynamically builds a positive and a negative cache to adapt features through the Tip-Adapter strategy Zhang et al.
+(2021).
+This method however requires to find specific hyperparameters for each dataset, as in the original Tip-Adapter approach.
+Pseudo-labels have been central to previous TTA methods.
 SHOT Liang et al.
-(2020) employs them with cross-entropy to regularize mutual information between inputs and predictions.
-The entire feature encoder is optimized in this case.
+(2020) uses them with cross-entropy to regularize mutual information, optimizing the entire feature encoder.
 CoTTA Wang et al.
-(2022) makes use of pseudo-labels in the context of a student-teacher model, where one model processes the original input and the other processes augmented versions of that input.
-A consistency loss is used to compare both models’ outputs.
+(2022) employs pseudo-labels in a student-teacher model with consistency loss between original and augmented inputs.
 PAD Wu et al.
-(2021) improves the quality of pseudo-labels by also augmenting inputs and using voting on the models’ predictions.
-Instead of simple class pseudo-labels, CLIPArTT improves the text supervision to broaden the likelihood of accurately predicting the right class.
-Some other methods utilize less conventional approaches for TTA.
+(2021) enhances pseudo-labels by augmenting inputs and voting on predictions.
+Instead of simple class pseudo-labels, our CLIPArTT method improves the text supervision to broaden the likelihood of accurately predicting the right class.
+Other methods take less conventional approaches to TTA.
 LAME Boudiaf et al.
-(2022) refines the classifier’s predictions in a transductive manner using the similarity between features through the Laplacian.
-A family of test-time training (TTT) methods have also gained popularity Sun et al.
+(2022) refines classifier predictions transductively using feature similarity through the Laplacian.
+Test-time training (TTT) methods Sun et al.
 (2020); Liu et al.
@@
 (2022); Vargas Hakim et al.
-(2023); Osowiechi et al.
-(2024), where a sub-branch is trained alongside the main network in an unsupervised way, which is later used to update the model.
-These methods require training the model from scratch on the source domain.
-As LAME, our method is related to the concept of Laplacian regularization, however we exploit this concept very differently.
-Whereas LAME employs it to refine predictions in a traditional transductive inference approach where the model is kept as is, our method exploits in a test-time adaptation loss enforcing the consistency between embeddings of related batch samples.
-Contrary to TTT methods, we do not train the model from scratch with any additional branch.
+(2023) train a sub-branch alongside the main network in an unsupervised manner to update the model, requiring training from scratch on the source domain.
+Like LAME, our method relates to Laplacian regularization but applies it differently.
+While LAME refines predictions without altering the model, our method uses it in a test-time adaptation loss to enforce consistency between embeddings of related batch samples.
+Unlike TTT methods, we do not require additional branches or training from scratch.
 Conformal Learning.
@@
 In a conformal prediction, a given level of certainty is assigned to a set $\mathcal{C}=\{c_{1},...,c_{K}\}$ with the $K$ most plausible classes that an input can belong to Angelopoulos and Bates (2021).
-We draw inspiration from this concept and build a conformal set of class predictions that can help adapting CLIP towards an accurate top--1 prediction.
+We draw inspiration from this concept and build a conformal set of class predictions that can help adapting CLIP towards an accurate top-1 prediction.
 Our technique, however, stands out by not relying on image transformations such as in Shu et al.
@@
 Building on the limitations of this approach, we then introduce our CLIPArTT method that leverages class uncertainty and the relationship between samples in a batch.
-Figure 1: CLIPArTT pipeline overview: 1) Computing predictions from Image-Text Similarity, 2) generating a new text prompt by filtering the top-$K$ class predictions, 3) employing the new prompt to obtain image-to-image and text-to-text similarity scores, used to build a pseudo-label for cross-entropy.
+Figure 1: CLIPArTT pipeline overview: 1) Computing predictions from Image-Text Similarity, 2) generating a new text prompt by filtering the top-$K$ class predictions, 3) with the new prompts, a pseudo-label Q is obtained by averaging the image-to-image and text-to-text similarity scores, while the prediction $\hat{\textbf{P}}$ is computed as the image-to-text similarity.
+Cross-entropy is then used as the TTA loss.
+Figure 2: Top: Example of similarity matrices ($\mathbf{S}^{v}$, $\mathbf{S}^{t}$) and CLIPArTT softmax probabilities ($\mathbf{Q}$) for a batch of 5 examples and using $K=5$ classes.
+Bottom: a) When using the identity matrix as pseudo-label for contrastive learning, the correct prediction is ambiguous, as the images are forced to both approaching and moving away from the right class.
+b) CLIPArTT uses soft pseudo-labels that smoothly guides the prediction towards the correct class by reducing the impact of ambiguities in the prompts.
 3.1 CLIP-based classification
@@
 For a new image $\mathbf{x}_{i}$, the probability of belonging to class $k$ is then estimated based on cosine similarity,
-$p_{ik}\,=\,\frac{\exp\big{(}\cos(\mathbf{z}^{v}_{i},\mathbf{z}^{t}_{k})/\tau\big{)}}{\sum_{j}\exp\big{(}\cos(\mathbf{z}^{v}_{i},\mathbf{z}^{t}_{j})/\tau\big{)}},\quad\cos(\mathbf{z},\mathbf{z}^{\prime})=\frac{{\mathbf{z}^{\top}}{\mathbf{z}^{\prime}}}{\|\mathbf{z}\|_{2}\!\cdot\!\|\mathbf{z}^{\prime}\|_{2}},$ (1)
+$p_{ik}\,=\,\frac{\exp\big{(}\!\cos(\mathbf{z}^{v}_{i},\mathbf{z}^{t}_{k})/\tau\big{)}}{\sum_{j}\exp\big{(}\!\cos(\mathbf{z}^{v}_{i},\mathbf{z}^{t}_{j})/\tau\big{)}},\ \ \cos(\mathbf{z},\mathbf{z}^{\prime})=\frac{{\mathbf{z}^{\top}}{\mathbf{z}^{\prime}}}{\|\mathbf{z}\|_{2}\!\cdot\!\|\mathbf{z}^{\prime}\|_{2}},$ (1)
 where $\tau$ is a suitable softmax temperature.
@@
 Original 88.74 97.79 61.68 80.92
-C-Average 59.22 82.43 29.43 46.61
+Corrupted (-C) 59.22 82.43 29.43 46.61
 Table 1: Accuracy (%) on CIFAR-10/100 and CIFAR-10/100-C datasets with Level 5 corruption for the top-1 or the top-3 predicted classes.
-Figure 2: Evolution of the accuracy after several iterations of adaptation at test-time for TENT.
+Refer to caption Refer to caption
+Figure 3: Evolution of CLIPArTT’s accuracy during test-time adaptation. Left: For different versions of CIFAR10. Right: Compared to TENT on CIFAR10-C.
 3.2 Our CLIPArTT method
-The proposed CLIPArTT method, illustrated in Figure 1, addresses the above-mentioned limitations using two key insights.
-The first insight, inspired by conformal learning, is that the correct class is often among the ones with the highest probabilities, although the model’s most confident prediction may not be always correct.
-This claim is supported by the results in Table 13 (Original), showing that the correct class is within the top-3 predictions 97.79% of the times for CIFAR10 (versus 88.74% within the top-1), and 80.92% of the times for CIFAR100 (versus 61.68% within the top-1).
+The proposed CLIPArTT method (see Figure 1), addresses the above-mentioned limitations using two key insights.
+The first insight, inspired by conformal learning, is that the correct class is often among the top most probable ones, although the model’s most confident prediction may not be always correct.
+This claim is supported by the results in Table 1 (first row: Original), showing that the correct class is within the top-3 predictions 97.79% of the times for CIFAR-10 (versus 88.74% within the top-1), and 80.92% of the times for CIFAR-100 (versus 61.68% within the top-1).
 A way to include information about multiple classes could, therefore, help adapting the model at test time.
 The second insight is that the similarity between the batch samples could be evaluated based on their visual and/or text embeddings.
-As we will show below, such similarities could be exploited in a strategy related to graph-Laplacian regularization.
+As we will show below, such similarities could be exploited in a strategy related to Stochastic Neighbor Embedding (SNE) and graph-Laplacian regularization.
 […]
-Transductive TTA  Next, we design a test-time adaptation loss that accounts for semantic relationships between batch samples.
+Transductive TTA.
+Next, we design a test-time adaptation loss that accounts for semantic relationships between batch samples.
 Let $\mathbf{Z}^{v}\in{}^{B\times D}$ and $\hat{\mathbf{Z}}^{t}\in{}^{B\times D}$ denote the normalized visual and instance-specific text embeddings of the samples within the test batch, respectively.
@@
 The latter captures common (or related) classes in the top-$k$ predictions of two samples, since it is computed using the instance-specific multi-class prompts.
+As illustrated in Figure 2 (top), a broad range of similarity values are obtained with this approach.
 We deploy these two pairwise similarity matrices to compute pseudo-labels as follows:
@@
 where the softmax operation is applied column-wise and the temperature $\tau=0.01$ in all our experiments.
-Let $\hat{\mathbf{P}}$ denotes the zero-shot prediction matrix using our instance-specific multi-class text prompts:
-$\hat{\mathbf{P}}=\text{softmax}\big{(}\mathbf{Z}^{v}{(\hat{\mathbf{Z}}^{t})^{\top}}/\tau\big{)},\quad\hat{p}_{ij}=\frac{\exp\big{(}\cos(\mathbf{z}^{v}_{i},\hat{\mathbf{z}}^{t}_{j})/\tau\big{)}}{\sum_{k}\exp\big{(}\cos(\mathbf{z}^{v}_{i},\hat{\mathbf{z}}^{t}_{k})/\tau\big{)}}$ (4)
+Let $\hat{\mathbf{P}}$ denote the zero-shot prediction matrix using our instance-specific multi-class text prompts:
+$\hat{\mathbf{P}}=\text{softmax}\big{(}\mathbf{Z}^{v}{(\hat{\mathbf{Z}}^{t})^{\top}}/\tau\big{)},\,\ \hat{p}_{ij}=\frac{\exp\big{(}\!\cos(\mathbf{z}^{v}_{i},\hat{\mathbf{z}}^{t}_{j})/\tau\big{)}}{\sum_{k}\exp\big{(}\!\cos(\mathbf{z}^{v}_{i},\hat{\mathbf{z}}^{t}_{k})/\tau\big{)}}$ (4)
 This matrix, along with the pairwise pseudo-labels we introduced in Eq. (3), yield our final TTA loss based on the cross-entropy:
 $\mathcal{L}_{\mathrm{TTA}}(\theta)=-\frac{1}{B}\sum_{i=1}^{B}\sum_{j=1}^{B}q_{ij}\log\hat{p}_{ij}.$ (5)
-Unlike recent approaches Shu et al.
-(2022), which adapt the CLIP model by learning a text prompt, we instead update the normalization-layer parameters of the visual encoder, which yields a light-weight, computationally efficient TTA method.
-3.3 Link to graph-Laplacian regularization
+Unlike recent approaches like TPT Shu et al.
+(2022), which adapt CLIP by learning a text prompt, we instead update the normalization-layer parameters of the visual encoder, which yields a light-weight, computationally efficient TTA method.
+We note that TPT requires to generate multiple augmentations during the forward pass, incurring a substantial overhead.
+The mechanism of CLIPArTT is illustrated in Fig.
+2.
+Our soft pseudo-label reduces the risk of increasing the ambiguity of the predictions.
+Using the identity matrix (i.e., hard pseudo-labels) leads to uncertain class assignments, as the ambiguity of the text prompts would both attract and repel the right class if it is present in the same text prompt for two different images.
+3.3 Link to existing techniques
 An important element of our TTA method is that it exploits the similarities between pairs of samples within the batch, in terms of the visual features and text prompts representing the top-$k$ classes.
-In this section, we draw a connection between the proposed method and graph-Laplacian regularization, well-known in the context semi-supervised learning techniques Chapelle et al.
+In this section, we draw a connection between the proposed method and two well-known techniques in machine learning: Stochastic Neighbor Embedding Hinton and Roweis (2002) and graph-Laplacian regularization Chapelle et al.
+(2006); Iscen et al.
+(2019).
+Stochastic Neighbor Embedding (SNE).
+This popular technique for dimensionality reduction estimates the local probability of a point $\mathbf{x}_{j}$ given a point $\mathbf{x}_{i}$ based on their Euclidean distance $d_{ij}=\|\mathbf{x}_{i}-\mathbf{x}_{j}\|_{2}$:
+$p_{ij}\,=\,\frac{\exp\big{(}-\!d^{2}_{ij}\big{)}}{\sum_{k\neq i}\exp\big{(}-\!d^{2}_{ik}\big{)}}$ (6)
+The goal is to find, for each $\mathbf{x}_{i}$, a low-dimensional embedding $\mathbf{y}_{i}$ such that the local probabilities $q_{ij}$ computed using Eq.
+(6) on these embeddings, is similar to $p_{ij}$.
+This is achieved by minimizing the row-wise KL divergence between distribution matrices $\mathbf{P}$ and $\mathbf{Q}$.
+Our CLIPArTT method can be linked to SNE since $\cos(\mathbf{x}_{i},\mathbf{x}_{j})=-\!d^{2}_{ij}+2$ when $\mathbf{x}_{i}$, $\mathbf{x}_{j}$ are unit normalized, and KL divergence between $\mathbf{P}$ and $\mathbf{Q}$ is equal to the cross-entropy between these matrices minus the entropy of $\mathbf{P}$.
+In summary, our TTA loss in Eq.
+(5) ensures that the inter-modality (text-to-image) similarities of batch samples are aligned with their intra-modality ones (text-to-text and image-to-image).
+Graph-Laplacian regularization  is commonly-used in the context semi-supervised learning techniques Chapelle et al.
 (2006) and label propagation Iscen et al.
 (2019).
-In our case, the Laplacian regularizer is computed over a set of $B$ nodes with predictions $\mathbf{Z}\in{}^{B\times D}$ as
+In our case, the Laplacian regularizer is computed over a set of $B$ nodes with predictions $\mathbf{Z}\in{}^{B\times D}$ as:
 $\mathcal{L}_{\mathrm{reg}}(\mathbf{Z})=\mathrm{trace}\big{(}{\mathbf{Z}^{\top}}\mathbf{L}_{\mathrm{W}}\mathbf{Z}\big{)}\,=\,\sum_{i,j}w_{ij}\|\mathbf{z}_{i}-\mathbf{z}_{j}\|_{2}^{2},$ (7)
@@
 $\mathcal{L}_{\mathrm{TTA}}$ $\,=\,-\frac{1}{B}\sum_{i=1}^{B}\sum_{j=1}^{B}q_{ij}\log\frac{\exp\big{(}{(\mathbf{z}^{v}_{i})^{\top}}\mathbf{z}^{t}_{j}/\tau\big{)}}{\sum_{k}\exp\big{(}{(\mathbf{z}^{v}_{i})^{\top}}\mathbf{z}^{t}_{k}/\tau\big{)}}$ (9)
-$=\frac{1}{\tau B}\sum_{i}\sum_{j}q_{ij}\Big{(}-\!{(\mathbf{z}^{v}_{i})^{\top}}\mathbf{z}^{t}_{j}\,+\,\tau\underbrace{\log\sum_{k}\exp\big{(}{(\mathbf{z}^{v}_{i})^{\top}}\mathbf{z}^{t}_{k}/\tau\big{)}}_{\text{LogSumExp (LSE)}}\Big{)}$ (9)
+$\,=\,\frac{1}{\tau B}\sum_{i}\sum_{j}q_{ij}\Big{(}-\!{(\mathbf{z}^{v}_{i})^{\top}}\mathbf{z}^{t}_{j}$
+$\qquad\qquad\quad\ +\,\tau\underbrace{\log\sum_{k}\exp\big{(}{(\mathbf{z}^{v}_{i})^{\top}}\mathbf{z}^{t}_{k}/\tau\big{)}}_{\text{LogSumExp (LSE)}}\Big{)}$ (10)
 Since the feature embeddings are normalized, we have $\|\mathbf{z}_{i}^{v}\|_{2}=\|\mathbf{z}_{j}^{t}\|_{2}=1$ and thus ${(\mathbf{z}^{v}_{i})^{\top}}\mathbf{z}^{t}_{j}=2-\|\mathbf{z}^{v}_{i}-\mathbf{z}^{t}_{j}\|_{2}^{2}$.
@@
 $LSE\,\approx\,\tfrac{1}{\tau}\max_{k}\,\{{(\mathbf{z}^{v}_{i})^{\top}}\mathbf{z}^{t}_{k}\}\,=\,\tfrac{1}{\tau}{(\mathbf{z}^{v}_{i})^{\top}}\mathbf{z}^{t}_{i}$ (12)
-The last equality comes from the hypothesis that, in a well trained model, the maximum similarity occurs between the image and embedding of a samples and its corresponding text embedding.
+The last equality comes from the hypothesis that, in a well trained model, the maximum similarity occurs between the image embedding of a sample and its corresponding text embedding.
 Our TTA loss can then be expressed as
-$\mathcal{L}_{\mathrm{TTA}}$ $\approx\frac{1}{\tau B}\Big{(}\sum_{i}\sum_{j}q_{ij}\|\mathbf{z}^{v}_{i}-\mathbf{z}^{t}_{j}\|_{2}^{2}\ +\,\ \sum_{i}{(\mathbf{z}^{v}_{i})^{\top}}\mathbf{z}^{t}_{i}\Big{)}\,+\,\text{const}$ (12)
-$=\frac{1}{\tau B}\mathrm{trace}\big{(}{(\mathbf{Z}^{v})^{\top}}\!(\mathbf{L}_{\mathrm{Q}}+\mathbf{I})\,\mathbf{Z}^{t}\big{)}\ +\,\text{const}$ (13)
+$\mathcal{L}_{\mathrm{TTA}}$ $\,\approx\,\frac{1}{\tau B}\Big{(}\sum_{i,j}q_{ij}\|\mathbf{z}^{v}_{i}-\mathbf{z}^{t}_{j}\|_{2}^{2}\,+\,\sum_{i}{(\mathbf{z}^{v}_{i})^{\top}}\mathbf{z}^{t}_{i}\Big{)}\,+\text{\emph{const}}$
+$\,=\,\frac{1}{\tau B}\mathrm{trace}\big{(}{(\mathbf{Z}^{v})^{\top}}\!(\mathbf{L}_{\mathrm{Q}}+\mathbf{I})\,\mathbf{Z}^{t}\big{)}\ +\text{\emph{const}}$ (13)
 where $\mathbf{I}$ is the identity matrix.
 The modified Laplacian $\mathbf{L}_{\mathrm{Q}}+\mathbf{I}$ enforces nodes with a high connection weight ($q_{ij}$) to have similar embeddings, while also avoiding embeddings to collapse into a single vector.
+∎
 4 Experimental Settings
-We evaluate CLIPArTT’s performance across diverse TTA datasets using established methodologies.
-These datasets simulate challenging domain shifts, offering insights into our approach’s efficacy.
-Additionally, we explore CLIPArTT’s adaptability on the original dataset through zero-shot test-time adaptation, crucial given the novel nature of the datasets.
-Our evaluation framework encompasses distinct categories, namely natural images, common corruptions, simulated shift and video shift, providing a comprehensive assessment of the model’s performance across diverse challenges.
-In our evaluation of natural images, we utilize CIFAR-10, CIFAR-10.1, and CIFAR-100 datasets, each comprising 10,000 images and featuring 10 and 100 classes, respectively.
-These datasets represent natural imagery and are novel to the model under scrutiny.
-Notably, CIFAR-10.1 introduces a natural domain shift from CIFAR-10, thereby enriching our assessment with varied and nuanced data distributions.
-Transitioning to our investigation of common corruptions, we turn to the CIFAR-10-C and CIFAR-100-C datasets Hendrycks and Dietterich (2019).
-These datasets offer a diverse range of 15 distinct corruptions, including elastic transform and impulse noise, among others.
-Each corruption is characterized by 5 severity levels, yielding a total of 75 unique testing scenarios per dataset.
-Within each severity level, there are 10,000 images, contributing to a comprehensive evaluation of the model’s robustness against a variety of corruption types and intensities.
-Furthermore, we examine the challenging VisDA-C dataset Peng et al.
-(2018), which represents two domain shifts: the simulated shift and video shift.
-The former comprises a set of 152,397 images rendered in 3D across 12 different classes.
-The later includes 72,372 YouTube video frames spanning the same categories.
-This dataset presents an important challenge, as it bridges the gap of the type of imagery that a model can be applied on.
+We evaluate CLIPArTT’s performance across a variety of TTA datasets, covering scenarios such as natural images, common corruptions, simulated images, video, and Domain Generalization benchmark.
+This comprehensive framework allows for a robust assessment of the model’s adaptability to various challenges, including domain shifts and corruptions.
+For a detailed description of the datasets, please refer to the supplementary materials.
 Test-time adaptation.
@@
 Throughout our experiments, a consistent batch size of 128 is utilized to maintain uniformity and enable effective comparisons across different scenarios.
-As in previous TTA works, a smaller learning rate of $10^{-4}$ is preferred to adapt on the 3D renderings split Vargas Hakim et al.
+As in previous TTA works, a smaller learning rate of $10^{-4}$ is preferred to adapt to the 3D renderings split Vargas Hakim et al.
 (2023), as it represents a more aggressive shift.
 Benchmarking.
-We compare the performance of CLIPArTT with state-of-the-art methods.
+We compare CLIPArTT with state-of-the-art methods.
 Specifically, we utilize adapted versions of TENT Wang et al.
@@
 (2022) tailored to CLIP.
-The adapted TENT now updates only the affine parameters in Layer Normalization layers through entropy minimization directly on the similarity between image and text features, diverging from its original implementation (i.e., using classifier’s logits).
+While the classifier logits are used in previous TTA research involving CNNs, we follow the standard practice and use the image-to-text similarity to obtain the final logits in CLIP.
+Only the visual encoder is optimized where needed.
+TENT now updates only the affine parameters in LN layers through entropy minimization directly on these logits.
 Notably, we employ 10 iterations for TENT adaptation, a choice informed by our findings, as evidenced in Fig 3.
-For VisDA-C, we also explore using only 1 iteration, as in the original paper.
-Similarly, we utilize the image-to-text similarity scores as logits for LAME, where the Laplacian regularizer is applied to the similarity between image features.
-Finally, we extend the evaluation to include TPT Shu et al.
-(2022), a recent adaptation method for CLIP.
-TPT heavily relies on image augmentations, thus we reduce batch-size of 32 due to its high memory requirement.
+In LAME, the Laplacian regularizer is applied on the similarity between image features (obtained from the visual encoder).
+Finally, we include CLIP-tailored methods TTDN Zhou et al.
+(2023), TPT Shu et al.
+(2022), TDA Karmanov et al.
+(2024) and CALIP Guo et al.
+(2023).
+Both TDA and CALIP require hyperparameter tuning for optimal performance.
+To ensure fairness in our benchmarking process, we opted to fix these hyperparameters as recommended for the ImageNet dataset, as suggested in their respective works.
+The batch-size for TPT is reduced to 32 due to its reliance on image augmentations and high memory requirements.
 $K$ = 1 $K$ = 3 $K$ = 4
-Original 89.80 $\pm$0.05 90.04 $\pm$0.13 90.41 $\pm$0.07
-CIFAR 10.1 85.37 $\pm$0.17 86.35 $\pm$0.27 86.07 $\pm$0.21
-C-Average 70.79 71.17 70.99
+CIFAR-10 89.80 $\pm$0.05 90.04 $\pm$0.13 90.41 $\pm$0.07
+CIFAR-10.1 85.37 $\pm$0.17 86.35 $\pm$0.27 86.07 $\pm$0.21
+CIFAR-10-C 70.79 $\pm$0.15 71.17 $\pm$0.16 70.99 $\pm$0.15
 Table 2: Accuracy (%) on CIFAR-10, CIFAR-10.1 and CIFAR-10-C datasets with Level 5 corruption for different number of $K$ selected classes to create pseudo-labels.
-$K$ = 3 $K$ = 4 $K$ = 5
-3D 84.31 84.24 87.24
-YT 82.80 83.17 83.46
-Table 3: Accuracy (%) on the VisDA-C 3D training split (3D) and validation split (YT) with ViT-B/32 on different numbers of selected classes ($K$) to create pseudo-labels.
+Iter = 1 Iter = 5 Iter = 10 Iter = 20
+CIFAR-10 89.59 $\pm$0.01 90.54 $\pm$0.09 90.04 $\pm$0.13 88.32 $\pm$0.12
+CIFAR-10.1 84.78 $\pm$0.02 86.67 $\pm$0.06 86.35 $\pm$0.27 84.33 $\pm$0.31
+CIFAR-10-C 62.30 $\pm$0.06 68.75 $\pm$0.12 71.17 $\pm$0.16 70.55 $\pm$0.24
+Table 3: Accuracy (%) on CIFAR-10, CIFAR-10.1 and CIFAR-10-C datasets with Level 5 corruption for different number of iterations to update the model at test-time.
 5 Results
@@
 5.1 Ablation Studies
-Selection of the number of classes for new prompts.
+Number of classes for new prompts.
 Determining the number of classes in building new prompts presents a critical aspect in our methodology.
-Table 13 illustrates various scenarios where $K\!=\!3$ emerges as a favorable choice, exhibiting a remarkable accuracy of $100\%$ across multiple instances of CIFAR-10 datasets.
+Table 1 illustrates various scenarios where $K\!=\!3$ emerges as a favorable choice, exhibiting a remarkable accuracy above $90\%$ across multiple instances of CIFAR-10-C datasets.
 This initial observation is further corroborated by the findings presented in Table 2.
@@
 Similarly, adopting a strategy akin to CIFAR-10, wherein $30\%$ of the classes are chosen, proves impractical due to resulting lengthy sentences that are impossible to tokenize.
-Examination of Table 4 reveals that the optimal choice varies depending on the distribution shift.
 Nonetheless, our analysis shows that $K\!=\!3$ consistently yields superior performance, particularly evident in the average results across CIFAR-100-C.
 Consequently, we maintain $K\!=\!3$ as the preferred configuration for all subsequent experiments on CIFAR datasets.
-A similar study was conducted for VisDA-C, in which $K\!=\!5$ (See Table 3) is chosen as the number of classes.
-$K$ = 1 $K$ = 3 $K$ = 5 $K$ = 7 $K$ = 10 $K$ = 20
-Original 69.00 $\pm$0.22 69.79 $\pm$0.04 69.68 $\pm$0.07 69.56 $\pm$0.02 69.78 $\pm$0.02 69.93 $\pm$0.08
-Gaussian Noise 26.05 $\pm$0.11 25.32 $\pm$0.14 24.69 $\pm$0.03 24.60 $\pm$0.03 24.70 $\pm$0.15 23.73 $\pm$0.07
-Shot noise 28.88 $\pm$0.11 27.90 $\pm$0.05 27.25 $\pm$0.26 26.75 $\pm$0.07 26.83 $\pm$0.26 25.73 $\pm$0.13
-Impulse Noise 24.04 $\pm$0.09 25.62 $\pm$0.09 25.12 $\pm$0.14 25.25 $\pm$0.14 24.95 $\pm$0.23 24.57 $\pm$0.12
-Defocus blur 49.03 $\pm$0.15 49.88 $\pm$0.23 49.75 $\pm$0.11 49.74 $\pm$0.25 49.62 $\pm$0.10 49.49 $\pm$0.07
-Glass blur 26.77 $\pm$0.14 27.89 $\pm$0.03 27.76 $\pm$0.23 27.28 $\pm$0.07 26.57 $\pm$0.07 25.52 $\pm$0.09
-Motion blur 46.50 $\pm$0.09 47.93 $\pm$0.14 47.48 $\pm$0.21 47.57 $\pm$0.10 47.53 $\pm$0.09 47.36 $\pm$0.09
-Zoom blur 52.08 $\pm$0.12 52.70 $\pm$0.06 52.22 $\pm$0.10 52.10 $\pm$0.24 52.62 $\pm$0.24 52.82 $\pm$0.09
-Snow 49.24 $\pm$0.07 49.72 $\pm$0.01 48.98 $\pm$0.08 48.87 $\pm$0.08 49.13 $\pm$0.08 49.54 $\pm$0.13
-Frost 49.91 $\pm$0.07 49.63 $\pm$0.12 48.43 $\pm$0.17 48.11 $\pm$0.06 48.72 $\pm$0.08 49.13 $\pm$0.10
-Fog 47.15 $\pm$0.04 48.77 $\pm$0.04 48.95 $\pm$0.18 48.78 $\pm$0.22 49.06 $\pm$0.05 48.74 $\pm$0.36
-Brightness 60.01 $\pm$0.08 61.27 $\pm$0.08 60.77 $\pm$0.16 60.89 $\pm$0.19 60.98 $\pm$0.18 61.03 $\pm$0.19
-Contrast 46.90 $\pm$0.21 48.55 $\pm$0.24 49.01 $\pm$0.14 49.07 $\pm$0.03 49.27 $\pm$0.09 49.08 $\pm$0.12
-Elastic transform 36.32 $\pm$0.10 37.45 $\pm$0.08 37.63 $\pm$0.12 37.31 $\pm$0.09 37.13 $\pm$0.16 36.94 $\pm$0.11
-Pixelate 32.52 $\pm$0.17 33.88 $\pm$0.14 34.40 $\pm$0.15 34.38 $\pm$0.16 34.65 $\pm$0.09 34.32 $\pm$0.02
-JPEG compression 35.81 $\pm$0.11 36.07 $\pm$0.32 35.77 $\pm$0.01 35.60 $\pm$0.10 35.63 $\pm$0.10 35.29 $\pm$0.14
-Average 40.75 41.51 41.21 41.09 41.16 40.89
-Table 4: Accuracy (%) on CIFAR-100 and CIFAR-100-C datasets with Level 5 corruption for different number of $K$ selected classes to create pseudo-labels.
-Selection of the number of iterations.
+Avg. Prompt Image Text Image + Text
+CIFAR-10 76.94 $\pm$0.41 90.18 $\pm$0.02 89.05 $\pm$0.14 90.04 $\pm$0.13
+CIFAR 10.1 67.83 $\pm$0.49 86.25 $\pm$0.37 84.85 $\pm$0.40 86.35 $\pm$0.27
+CIFAR-10-C 43.08 $\pm$0.08 70.98 $\pm$0.15 70.03 $\pm$0.21 71.17 $\pm$0.16
+Table 4: Accuracy (%) on CIFAR-10, CIFAR-10.1 and CIFAR-10-C datasets with different targets.
+CLIP BS = 16 BS = 32 BS = 64 BS = 128
+CIFAR-10 88.74 85.89 $\pm$0.19 88.25 $\pm$0.15 89.48 $\pm$0.15 90.04 $\pm$0.13
+CIFAR-10.1 83.25 81.55 $\pm$0.53 84.00 $\pm$0.31 85.40 $\pm$0.08 86.35 $\pm$0.27
+CIFAR-10-C 59.22 64.72 $\pm$0.23 67.70 $\pm$0.23 69.82 $\pm$0.20 71.17 $\pm$0.16
+Table 5: Accuracy (%) on CIFAR-10, CIFAR-10.1 and CIFAR-10-C datasets with Level 5 corruption for different number of batch-size.
+CIFAR-100-C ImageNet-C
+CLIP TENT TDA CLIPArTT CLIP TENT TDA CLIPArTT
+Gaussian Noise 14.80 14.38 $\pm$0.14 08.20 $\pm$0.35 25.32 $\pm$0.14 12.27 12.28 $\pm$0.05 11.54 $\pm$0.05 20.18 $\pm$0.61
+Shot noise 16.03 17.34 $\pm$0.27 09.58 $\pm$0.43 27.90 $\pm$0.05 12.20 08.46 $\pm$0.08 12.13 $\pm$0.13 20.94 $\pm$0.08
+Impulse Noise 13.85 10.03 $\pm$0.13 07.63 $\pm$0.19 25.62 $\pm$0.09 12.89 15.71 $\pm$0.04 12.12 $\pm$0.08 19.95 $\pm$0.02
+Defocus blur 36.74 49.05 $\pm$0.07 25.59 $\pm$0.41 49.88 $\pm$0.23 21.60 20.61 $\pm$7.05 21.39 $\pm$0.08 24.51 $\pm$0.13
+Glass blur 14.19 03.71 $\pm$0.07 09.83 $\pm$0.56 27.89 $\pm$0.03 10.84 17.20 $\pm$0.08 10.74 $\pm$0.06 19.51 $\pm$0.02
+Motion blur 36.14 46.62 $\pm$0.27 28.92 $\pm$0.18 47.93 $\pm$0.14 17.85 24.60 $\pm$0.04 18.45 $\pm$0.06 25.80 $\pm$1.16
+Zoom blur 40.24 51.84 $\pm$0.15 31.08 $\pm$0.36 52.70 $\pm$0.06 16.38 21.13 $\pm$0.08 16.83 $\pm$0.03 24.00 $\pm$0.06
+Snow 38.95 46.71 $\pm$0.21 32.94 $\pm$0.12 49.72 $\pm$0.01 21.91 24.19 $\pm$0.09 22.97 $\pm$0.05 27.26 $\pm$0.06
+Frost 40.56 44.90 $\pm$0.27 34.84 $\pm$0.25 49.63 $\pm$0.12 23.33 23.34 $\pm$0.06 24.68 $\pm$0.06 26.56 $\pm$0.08
+Fog 38.00 47.31 $\pm$0.04 31.13 $\pm$0.15 48.77 $\pm$0.04 26.39 28.58 $\pm$0.04 27.72 $\pm$0.05 34.68 $\pm$0.02
+Brightness 48.18 60.58 $\pm$0.18 42.36 $\pm$0.10 61.27 $\pm$0.08 45.87 47.16 $\pm$0.01 48.06 $\pm$0.04 47.21 $\pm$0.07
+Contrast 29.53 45.90 $\pm$0.11 18.03 $\pm$0.07 48.55 $\pm$0.24 16.16 21.62 $\pm$0.01 16.09 $\pm$0.05 23.38 $\pm$0.12
+Elastic transform 26.33 33.09 $\pm$0.08 18.88 $\pm$0.24 37.45 $\pm$0.08 16.04 17.17 $\pm$0.01 17.75 $\pm$0.01 23.95 $\pm$0.04
+Pixelate 21.98 26.47 $\pm$0.09 14.59 $\pm$0.30 33.88 $\pm$0.14 28.00 33.56 $\pm$1.71 30.03 $\pm$0.05 34.12 $\pm$0.09
+JPEG compression 25.91 29.89 $\pm$0.07 17.56 $\pm$0.11 36.07 $\pm$0.32 27.44 31.89 $\pm$0.02 29.36 $\pm$0.01 32.23 $\pm$0.12
+Average 29.43 35.19 22.08 41.51 20.61 23.17 21.32 26.95
+Table 6: Accuracy (%) on CIFAR-100-C and ImageNet-C datasets with ViT-B/32 as visual encoder.
+[…]
+Number of iterations.
 In line with our approach for TENT, we investigate optimal number of iterations.
@@
 As a result, 10 iterations are used for all forthcoming experiments.
-Iter = 1 Iter = 5 Iter = 10 Iter = 20
-Original 89.59 $\pm$0.01 90.54 $\pm$0.09 90.04 $\pm$0.13 88.32 $\pm$0.12
-CIFAR 10.1 84.78 $\pm$0.02 86.67 $\pm$0.06 86.35 $\pm$0.27 84.33 $\pm$0.31
-Gaussian Noise 39.75 $\pm$0.04 53.79 $\pm$0.08 59.90 $\pm$0.36 59.33 $\pm$0.20
-Shot noise 43.80 $\pm$0.04 57.16 $\pm$0.24 62.77 $\pm$0.07 63.08 $\pm$0.43
-Impulse Noise 45.19 $\pm$0.07 52.44 $\pm$0.14 56.02 $\pm$0.16 56.73 $\pm$0.01
-Defocus blur 72.93 $\pm$0.07 76.36 $\pm$0.10 76.74 $\pm$0.05 75.33 $\pm$0.13
-Glass blur 46.61 $\pm$0.06 57.45 $\pm$0.13 61.77 $\pm$0.16 62.01 $\pm$0.27
-Motion blur 67.89 $\pm$0.03 74.34 $\pm$0.14 76.01 $\pm$0.19 75.94 $\pm$0.28
-Zoom blur 73.24 $\pm$0.05 77.03 $\pm$0.07 77.40 $\pm$0.20 75.42 $\pm$0.13
-Snow 73.81 $\pm$0.07 76.51 $\pm$0.08 77.29 $\pm$0.16 76.18 $\pm$0.21
-Frost 74.80 $\pm$0.04 78.13 $\pm$0.12 79.20 $\pm$0.08 77.44 $\pm$0.30
-Fog 69.81 $\pm$0.06 74.16 $\pm$0.09 75.74 $\pm$0.14 74.66 $\pm$0.02
-Brightness 84.16 $\pm$0.01 86.61 $\pm$0.10 86.59 $\pm$0.16 85.14 $\pm$0.42
-Contrast 67.75 $\pm$0.03 74.85 $\pm$0.04 77.82 $\pm$0.14 77.75 $\pm$0.11
-Elastic transform 63.15 $\pm$0.08 68.53 $\pm$0.19 70.20 $\pm$0.01 68.48 $\pm$0.24
-Pixelate 54.20 $\pm$0.02 61.87 $\pm$0.04 66.52 $\pm$0.13 67.13 $\pm$0.15
-JPEG compression 57.46 $\pm$0.09 62.00 $\pm$0.13 63.51 $\pm$0.14 63.64 $\pm$0.20
-Average 62.30 68.75 71.17 70.55
-Table 5: Accuracy (%) on CIFAR-10, CIFAR-10.1 and CIFAR-10-C datasets with Level 5 corruption for different number of iterations to update the model at test-time.
 Pseudo-label selection for adaptation.
@@
 Conversely, utilizing only the image-to-image similarity mainly improves for the original CIFAR-10 dataset and few corruption types.
-Image Text Image + Text
-Original 90.18 $\pm$0.02 89.05 $\pm$0.14 90.04 $\pm$0.13
-CIFAR 10.1 86.25 $\pm$0.37 84.85 $\pm$0.40 86.35 $\pm$0.27
-Gaussian Noise 59.05 $\pm$0.30 59.29 $\pm$0.27 59.90 $\pm$0.36
-Shot noise 62.11 $\pm$0.18 61.89 $\pm$0.31 62.77 $\pm$0.07
-Impulse Noise 55.43 $\pm$0.12 55.48 $\pm$0.10 56.02 $\pm$0.16
-Defocus blur 76.88 $\pm$0.09 76.25 $\pm$0.10 76.74 $\pm$0.05
-Glass blur 61.56 $\pm$0.03 61.28 $\pm$0.33 61.77 $\pm$0.16
-Motion blur 76.32 $\pm$0.15 75.37 $\pm$0.27 76.01 $\pm$0.19
-Zoom blur 77.66 $\pm$0.09 76.29 $\pm$0.11 77.40 $\pm$0.20
-Snow 77.28 $\pm$0.08 76.03 $\pm$0.16 77.29 $\pm$0.16
-Frost 78.80 $\pm$0.18 78.05 $\pm$0.21 79.20 $\pm$0.08
-Fog 75.69 $\pm$0.17 73.70 $\pm$0.15 75.74 $\pm$0.14
-Brightness 86.62 $\pm$0.20 84.69 $\pm$0.04 86.59 $\pm$0.16
-Contrast 77.43 $\pm$0.12 74.52 $\pm$0.02 77.82 $\pm$0.14
-Elastic transform 69.63 $\pm$0.02 69.33 $\pm$0.20 70.20 $\pm$0.01
-Pixelate 66.33 $\pm$0.16 64.86 $\pm$0.29 66.52 $\pm$0.13
-JPEG compression 63.92 $\pm$0.13 63.44 $\pm$0.20 63.51 $\pm$0.14
-Average 70.98 70.03 71.17
-Table 6: Accuracy (%) on CIFAR-10, CIFAR-10.1 and CIFAR-10-C datasets with different targets.
-CLIP BS = 8 BS = 16 BS = 32 BS = 64 BS = 128
-Original 88.74 82.50 $\pm$0.13 85.89 $\pm$0.19 88.25 $\pm$0.15 89.48 $\pm$0.15 90.04 $\pm$0.13
-CIFAR 10.1 83.25 77.2 $\pm$0.92 81.55 $\pm$0.53 84.00 $\pm$0.31 85.40 $\pm$0.08 86.35 $\pm$0.27
-C-Average 59.22 61.10 64.72 67.70 69.82 71.17
-Table 7: Accuracy (%) on CIFAR-10, CIFAR-10.1 and CIFAR-10-C datasets with Level 5 corruption for different number of batch-size.
 Performance of the model for different batch-sizes.
@@
 However, beyond a batch size of 8, no further performance gains are observed.
-This phenomenon can be attributed to the fact that, with smaller batch sizes, our method has the potential to introduce uncertainty by introducing multiple classes, thereby diminishing performance, especially when the model is already confident.
+This phenomenon can be attributed to the fact that, with smaller batch sizes, our method can potentially introduce uncertainty by using multiple classes, thereby diminishing performance, especially when the model is already confident.
 In subsequent experiments, we maintain a batch size of 128, consistent with prevailing practices in the state-of-the-art methodologies.
-5.2 Experiments
-CLIP LAME TENT TPT (BS=32) CLIPArTT
-Original 88.74 89.36 $\pm$0.06 91.69 $\pm$0.10 88.06 $\pm$0.06 90.04 $\pm$0.13
-CIFAR 10.1 83.25 81.22 $\pm$0.33 87.60 $\pm$0.45 81.80 $\pm$0.27 86.35 $\pm$0.27
-Gaussian Noise 35.27 24.71 $\pm$0.11 41.27 $\pm$0.27 33.90 $\pm$0.08 59.90 $\pm$0.36
-Shot noise 39.67 27.44 $\pm$0.09 47.20 $\pm$0.23 38.20 $\pm$0.02 62.77 $\pm$0.07
-Impulse Noise 42.61 31.38 $\pm$0.15 48.58 $\pm$0.31 37.66 $\pm$0.20 56.02 $\pm$0.16
-Defocus blur 69.76 62.45 $\pm$0.44 77.12 $\pm$0.16 67.83 $\pm$0.28 76.74 $\pm$0.05
-Glass blur 42.40 29.96 $\pm$0.06 52.65 $\pm$0.30 38.81 $\pm$0.12 61.77 $\pm$0.16
-Motion blur 63.97 54.00 $\pm$0.36 71.25 $\pm$0.09 63.39 $\pm$0.13 76.01 $\pm$0.19
-Zoom blur 69.83 61.97 $\pm$0.36 76.20 $\pm$0.19 68.95 $\pm$0.16 77.40 $\pm$0.20
-Snow 71.78 64.61 $\pm$0.48 78.29 $\pm$0.20 70.16 $\pm$0.10 77.29 $\pm$0.16
-Frost 72.86 65.17 $\pm$0.17 79.84 $\pm$0.09 72.39 $\pm$0.22 79.20 $\pm$0.08
-Fog 67.04 59.13 $\pm$0.49 77.39 $\pm$0.01 64.31 $\pm$0.28 75.74 $\pm$0.14
-Brightness 81.87 80.05 $\pm$0.23 87.78 $\pm$0.03 81.30 $\pm$0.18 86.59 $\pm$0.16
-Contrast 64.37 56.91 $\pm$0.37 79.47 $\pm$0.11 62.26 $\pm$0.31 77.82 $\pm$0.14
-Elastic transform 60.83 53.89 $\pm$0.20 70.00 $\pm$0.25 56.43 $\pm$0.27 70.20 $\pm$0.01
-Pixelate 50.53 39.67 $\pm$0.34 63.74 $\pm$0.18 42.80 $\pm$0.40 66.52 $\pm$0.13
-JPEG compression 55.48 47.24 $\pm$0.14 62.64 $\pm$0.14 53.67 $\pm$0.25 63.51 $\pm$0.14
-Average 59.22 50.57 67.56 56.80 71.17
+5.2 Comparison on different datasets
+In standard TTA, the concept of domain shift is strictly related to the source dataset used for pretraining.
+In the context of CLIP, the presence of domain shifts is less evident, as the source data contained an enormous amount of images that likely include different domains.
+For this reason, we categorize our experiments per dataset type.
+Shift CLIP LAME TENT TPT (BS=32) TTDN TDA CALIP CLIPArTT
+CIFAR-10 88.74 89.36 $\pm$0.06 91.69 $\pm$0.10 88.06 $\pm$0.06 89.06 $\pm$0.02 84.09 $\pm$0.04 86.79  $\pm$0.01 90.04 $\pm$0.13
+CIFAR-10.1 83.25 81.22 $\pm$0.33 87.60 $\pm$0.45 81.80 $\pm$0.27 83.77 $\pm$0.06 78.98 $\pm$0.37 81.02  $\pm$0.03 86.35 $\pm$0.27
+CIFAR-100 61.68 58.27 $\pm$0.17 69.74 $\pm$0.16 63.78 $\pm$0.28 64.10 $\pm$0.05 60.32 $\pm$0.06 61.94 $\pm$0.01 69.79 $\pm$0.04
+ImageNet 61.81 61.17 $\pm$0.02 62.13 $\pm$0.02 60.74 $\pm$0.16 61.85 $\pm$0.02 64.49 $\pm$0.01 58.50 $\pm$0.00 61.39 $\pm$0.09
+VisDA-C (YT) 84.45 84.72 $\pm$0.01 84.41 $\pm$0.02 82.95 $\pm$0.01 83.29 $\pm$0.03 84.47 $\pm$0.08 85.02 $\pm$0.00 83.46 $\pm$0.03
+PACS 93.65 93.68 $\pm$0.04 93.81 $\pm$0.03 93.23 $\pm$0.12 94.77 $\pm$0.07 92.94 $\pm$1.08 94.11 $\pm$0.00 93.95 $\pm$0.06
+Office Home 77.53 75.46 $\pm$0.12 77.68 $\pm$0.06 77.20 $\pm$0.24 79.76 $\pm$0.04 77.70 $\pm$0.13 78.75 $\pm$0.01 77.56 $\pm$0.09
+CIFAR-10-C 59.22 50.57 $\pm$0.30 67.56 $\pm$0.19 56.80 $\pm$0.22 60.13 $\pm$0.07 48.00 $\pm$0.68 56.58 $\pm$0.0 71.17 $\pm$0.16
+CIFAR-100-C 29.43 26.23 $\pm$0.12 35.19 $\pm$0.16 30.46 $\pm$0.14 31.90 $\pm$0.04 22.08 $\pm$0.29 29.92 $\pm$0.01 41.51 $\pm$0.15
+ImageNet-C 20.61 20.02 $\pm$0.16 23.17 $\pm$0.54 21.10 $\pm$0.06 21.28 $\pm$0.02 21.32 $\pm$0.06 21.00 $\pm$0.00 26.95 $\pm$0.12
+VisDA-C (3D) 84.51 84.75 $\pm$0.17 83.85 $\pm$0.03 78.45 $\pm$0.05 80.45 $\pm$0.01 84.22 $\pm$0.10 83.81 $\pm$0.08 87.24 $\pm$0.04
 Table 7: Accuracy (%) on CIFAR-10, CIFAR-10.1 and CIFAR-10-C datasets with ViT-B/32 as visual encoder.
-Performance analysis under natural or no domain shift.
-In the different Tables 8, 9, 10, and 11, CLIPArTT consistently enhances accuracy across CIFAR-10, CIFAR10.1, and CIFAR-100 datasets compared to the baseline (+2%, +3%, +8% respectively with ViT-B/32).
-This improvement suggests that even datasets devoid of corruption, which may not have been seen by the model during training, can benefit from adaptation.
+The Shift column denotes a domain shift from standard images recognized by CLIP: corruption in CIFAR-10-C, CIFAR-100-C and ImageNet-C datasets, synthetic images in VisDA-C (3D).
+Natural images.
+In the different Tables 7 and 8, CLIPArTT consistently enhances accuracy across CIFAR-10, CIFAR10.1, and CIFAR-100 datasets compared to the baseline (+2%, +3%, +8% respectively with ViT-B/32).
+This improvement suggests that uncorrupted datasets, which may not have been seen by the model during training, can benefit from adaptation.
 Consequently, a zero-shot model can leverage adaptation at test-time, relying solely on the model’s predictions.
-However, while CLIPArTT outperforms the baseline, as well as LAME and TPT with this fixed version, it is worth noting that TENT may yield superior results depending on the visual encoder and the number of classes.
+However, while CLIPArTT outperforms the baseline, as well as LAME, TPT and TTDN, it is worth noting that TENT may yield superior results depending on the visual encoder and the number of classes.
 Entropy minimization used solely on confident results proves to be also effective.
-Interestingly, the corruptions in which TENT achieves a better performance tend to already have higher accuracy values.
-CLIP LAME TENT TPT (BS=32) CLIPArTT
-Original 61.68 58.27 $\pm$0.17 69.74 $\pm$0.16 63.78 $\pm$0.28 69.79 $\pm$0.04
-Gaussian Noise 14.80 12.72 $\pm$0.04 14.38 $\pm$0.14 14.03 $\pm$0.10 25.32 $\pm$0.14
-Shot noise 16.03 13.78 $\pm$0.08 17.34 $\pm$0.27 15.25 $\pm$0.17 27.90 $\pm$0.05
-Impulse Noise 13.85 07.82 $\pm$0.14 10.03 $\pm$0.13 13.01 $\pm$0.13 25.62 $\pm$0.09
-Defocus blur 36.74 33.38 $\pm$0.11 49.05 $\pm$0.07 37.60 $\pm$0.17 49.88 $\pm$0.23
-Glass blur 14.19 09.00 $\pm$0.05 03.71 $\pm$0.07 16.41 $\pm$0.02 27.89 $\pm$0.03
-Motion blur 36.14 32.79 $\pm$0.13 46.62 $\pm$0.27 37.52 $\pm$0.23 47.93 $\pm$0.14
-Zoom blur 40.24 37.57 $\pm$0.15 51.84 $\pm$0.15 42.99 $\pm$0.11 52.70 $\pm$0.06
-Snow 38.95 35.49 $\pm$0.18 46.71 $\pm$0.21 42.35 $\pm$0.13 49.72 $\pm$0.01
-Frost 40.56 37.22 $\pm$0.21 44.90 $\pm$0.27 43.31 $\pm$0.14 49.63 $\pm$0.12
-Fog 38.00 35.94 $\pm$0.09 47.31 $\pm$0.04 38.81 $\pm$0.17 48.77 $\pm$0.04
-Brightness 48.18 44.93 $\pm$0.08 60.58 $\pm$0.18 50.23 $\pm$0.11 61.27 $\pm$0.08
-Contrast 29.53 27.52 $\pm$0.06 45.90 $\pm$0.11 28.09 $\pm$0.09 48.55 $\pm$0.24
-Elastic transform 26.33 24.01 $\pm$0.02 33.09 $\pm$0.08 28.12 $\pm$0.15 37.45 $\pm$0.08
-Pixelate 21.98 19.55 $\pm$0.13 26.47 $\pm$0.09 20.43 $\pm$0.14 33.88 $\pm$0.14
-JPEG compression 25.91 21.77 $\pm$0.14 29.89 $\pm$0.07 28.82 $\pm$0.09 36.07 $\pm$0.32
-Average 29.43 26.23 35.19 30.46 41.51
-Table 9: Accuracy (%) on CIFAR-100 and CIFAR-100-C datasets with ViT-B/32 as visual encoder.
-Performance analysis under common corruptions.
-In the various referenced tables, including Tables 8, 9, 10, and 11, CLIPArTT consistently outperforms the baseline, TPT, and LAME across all visual encoders, resulting in enhancements of up to 13% on average.
+Interestingly, the corruptions in which TENT achieves a better performance tend to already have higher accuracy.
+For a larger dataset such as ImageNet, most methods perform similarly, with the exception of TDA and CALIP, which improve CLIP performance by over 2%.
+Notably, both TDA and CALIP require hyperparameter tuning, and we used the recommended settings optimized for ImageNet.
+Varied styles and textures.
+Next we compared methods on two datasets, PACS and OfficeHome, which contain images from very different domains and are often used as benchmark in domain generalization.
+Once again, these domains do not constitute real distribution shifts for CLIP which was trained with a broad range of datasets.
+As can be seen, the best performance is achieved by methods such as TTDN for this setting.
+Nevertheless, it is important to highlight CLIPArTT’s robustness, as it consistently maintains or improves model performance after adaptation.
+For instance, on the PACS dataset, CLIPArTT achieves an accuracy of 93.95%, compared to 93.65% for the original CLIP.
+Backbone CLIP TENT CLIPArTT
+CIFAR-10.1 ViT-B/16 84.00 88.52 $\pm$0.33 88.72 $\pm$0.33
+ViT-L/14 91.20 92.22 $\pm$0.25 91.02 $\pm$0.02
+CIFAR-10-C ViT-B/16 60.15 68.00 $\pm$0.18 73.22 $\pm$0.17
+ViT-L/14 76.04 79.18 $\pm$0.10 78.06 $\pm$0.15
+CIFAR-100-C ViT-B/16 32.01 37.90 $\pm$0.18 40.08 $\pm$0.18
+ViT-L/14 44.59 50.14 $\pm$0.15 52.52 $\pm$0.18
+Table 8: Accuracy (%) on CIFAR-10, CIFAR-10.1, CIFAR-10-C, CIFAR-100 and CIFAR-100-C datasets with ViT-B/16 and ViT-L/14 as visual encoders.
+Common corruptions.
+In the various referenced tables, including Tables 7 and 8, CLIPArTT consistently outperforms the most benchmarks across all visual encoders, resulting in enhancements of up to 13% on average.
 A more detailed examination reveals that CLIPArTT frequently surpasses TENT, particularly in instances where the baseline performs poorly.
@@
 For instance, in Table 6, TENT experiences a 3% decrease compared to the baseline under Impulse Noise, while CLIPArTT exhibits a 12% improvement.
+In Figure 3 (right), it is evident that CLIPArTT achieves better performance more rapidly across varying numbers of iterations.
 Conversely, when the baseline is confident, CLIPArTT consistently maintains a high performance.
-As an illustration, Table 8 shows that while CLIP achieves an accuracy of 81.87% under Brightness corruption, both TENT and CLIPArTT contribute to improvements of 7% and 6%, respectively.
-CLIP TENT TPT (BS=32) CLIPArTT
-Original ViT-B/16 89.25 92.75 $\pm$0.17 89.80 $\pm$0.05 92.61 $\pm$0.05
-ViT-L/14 95.36 96.13 $\pm$0.06 95.18 $\pm$0.02 95.16 $\pm$0.03
-CIFAR-10.1 ViT-B/16 84.00 88.52 $\pm$0.33 83.75 $\pm$0.21 88.72 $\pm$0.33
-ViT-L/14 91.20 92.22 $\pm$0.25 91.32 $\pm$0.12 91.02 $\pm$0.02
-C-Average ViT-B/16 60.15 68.00 59.75 73.22
-ViT-L/14 76.04 79.18 75.01 78.06
-Table 10: Accuracy (%) on CIFAR-10, CIFAR-10.1 and CIFAR-10-C datasets with ViT-B/16 and ViT-L/14 as visual encoders.
-CLIP TENT TPT (BS=32) CLIPArTT
-Original ViT-B/16 64.76 71.73 $\pm$0.14 67.15 $\pm$0.23 71.34 $\pm$0.07
-ViT-L/14 73.28 78.03 $\pm$0.08 76.85 $\pm$0.06 79.42 $\pm$0.08
-C-Average ViT-B/16 32.01 37.90 33.73 40.08
-ViT-L/14 44.59 50.14 47.58 52.52
-Table 11: Accuracy (%) on CIFAR-100 and CIFAR-100-C datasets with ViT-B/16 and ViT-L/14 as visual encoders.
-CLIP TENT TENT-10 LAME TPT (BS=32) CLIPArTT
-3D ViT-B/16 87.04 87.15 $\pm$0.21 88.08 $\pm$0.05 85.83 $\pm$0.06 83.85 $\pm$0.06 88.01 $\pm$0.47
-ViT-B/32 84.51 84.75 $\pm$0.17 85.82 $\pm$0.02 83.85 $\pm$0.03 78.45 $\pm$0.05 87.24 $\pm$0.04
-ViT-L/14 90.62 90.93 $\pm$0.01 91.33 $\pm$0.01 90.15 $\pm$0.07 91.07 $\pm$0.05 91.18 $\pm$0.01
-YT ViT-B/16 86.71 86.96 $\pm$0.02 87.44 $\pm$0.02 86.24 $\pm$0.0 85.69 $\pm$0.05 86.22 $\pm$0.01
-ViT-B/32 84.45 84.72 $\pm$0.01 85.46 $\pm$0.02 84.41 $\pm$0.02 82.95 $\pm$0.01 83.46 $\pm$0.03
-ViT-L/14 84.93 85.38 $\pm$0.01 86.65 $\pm$0.05 85.08 $\pm$0.04 85.54 $\pm$0.38 85.55 $\pm$0.01
-Table 12: Accuracy (%) on the VisDA-C training split (3D) and validation split (YT) with the three different feature extractors.
-Performance analysis under simulated and video shifts.
-Results on the 3D and YT splits of VisDA-C are shown in Table 12.
-With respect to pure CLIP, our method achieves an important improvement of accuracy with the three different image encoders.
-CLIPArTT also achieves a competitive advantage against TENT, LAME and TPT.
-We observed however that using TENT with 10 iterations (here named TENT-10) is a strong contender, achieving practically the same score on 3D images and outperforming on YouTube frames by a small margin.
-This can mean that TENT’s entropy minimization can achieve faster improvement with certain domain shifts.
+Finally, our method performs effectively on larger datasets with a greater number of classes, consistently improving upon the baseline by 6% on ImageNet-C.
+Simulated images and video.
+The performance of CLIPArTT on the YT dataset is suboptimal, a trend that can be observed in scenarios where the CLIP model already exhibits high confidence.
+However, our observations indicate that on the 3D dataset, CLIPArTT demonstrates a notable competitive advantage over other Test-Time Adaptation (TTA) methods and the baseline, achieving an improvement of nearly 3%.
+Additionally, while TENT – often considered one of the more robust methods in handling various corruptions – results in a performance decline relative to the baseline, CLIPArTT maintains its effectiveness.
+5.3 Limitations
+Our results indicate that CLIPArTT performs particularly well under significant domain shifts such as corruptions, though it is slightly less effective on natural images.
+While the performance of CLIPArTT in these cases is not as strong as that of TENT, it remains stable without deteriorating CLIP’s capabilities, whereas the performance of TENT degrades under severe domain shifts.
+It is important to mention that any dataset not included in CLIP’s pre-training can be considered a potential domain shift.
+However, natural images are more likely to have been seen by CLIP during pre-training, making adaptation for them prone to overfitting.
+This may explain why CLIPArTT excels in more challenging scenarios.
+To further explore the potential limitations of our method, we have included two additional scenarios in the supplementary material: a) evaluation in a highly imbalanced setting, where only $C$ randomly chosen classes are present in a batch at random, and b) open-set classification, where out-of-distribution images (i.e., from classes not present in the text prompts) are introduced into the batches.
 6 Conclusions
-In this study, we introduce CLIPArTT, the first comprehensive Test-Time Adaptation framework designed specifically for Vision-Language Models.
+We introduce CLIPArTT, a novel Test-Time Adaptation framework designed specifically for VLMs.
 By leveraging the model’s predictions as new pseudo-labels, we effectively minimize cross-entropy loss at test-time, thereby enhancing model performance even within a zero-shot setting.
 To ensure fair comparisons, we also incorporate Test-Time Adaptation models into this new benchmark for VLMs.
-A comprehensive ablation study is conducted to determine optimal hyperparameters and gain deeper insights into the various model configurations.
+A comprehensive ablation study determined optimal hyperparameters and helped gain deeper insights into the various model configurations.
 Our experimental results demonstrate that CLIPArTT achieves highly competitive performance across TTA datasets, surpassing state-of-the-art approaches.
 While TENT remains a strong competitor, our model offers enhanced versatility by effectively addressing both natural and severe domain shifts, thus exhibiting robustness across various scenarios.
-Exploring the potential of text prompts in classification represents a promising avenue for future research.
+Exploring the potential of text prompts in classification is a promising avenue for future research.
 Investigating alternative methods to fine-tune these text prompts could yield valuable insights.
@@
 CLIPArTT: Light-weight Adaptation of CLIP to New Domains at Test Time – Supplementary Material
-Appendix A Unsupervised clustering
+Appendix A Dataset Details
+We evaluate CLIPArTT’s performance across diverse TTA datasets using established methodologies.
+These datasets simulate challenging scenarios, offering insights into our approach’s efficacy.
+Additionally, we explore CLIPArTT’s adaptability on other datasets through zero-shot test-time adaptation.
+Our evaluation framework encompasses natural images, varied styles and textures images, common corruptions, simulated images, and video providing a comprehensive assessment of the model’s performance across diverse challenges.
+In our evaluation of natural images (also known as zero-shot scenario), we utilize CIFAR-10, CIFAR-10.1, and CIFAR-100 datasets, each comprising 10,000 images and featuring 10 and 100 classes, respectively.
+These datasets represent natural imagery and are novel to the model under scrutiny.
+Notably, CIFAR-10.1 introduces a natural domain shift from CIFAR-10, thereby enriching our assessment with varied and nuanced data distributions.
+We also evaluate our method on ImageNet and extend our investigation on two datasets mostly used in the field of domain generalization: PACS Li et al.
+(2017) and OfficeHome Venkateswara et al.
+(2017) datasets, instrumental in understanding texture and style variations.
+The PACS dataset consists of 9,991 images across four domains (Art, Cartoons, Photos, Sketches) and seven classes.
+Lastly, the OfficeHome dataset includes 15,588 images across four domains (Art, Clipart, Product, Real) and 65 classes.
+Evaluating across these distinct scenarios showcases the generalizability of our method.
+Transitioning to our investigation of common corruptions, we turn to the CIFAR-10-C and CIFAR-100-C Hendrycks and Dietterich (2019).
+These datasets offer a diverse range of 15 distinct corruptions, including elastic transform and impulse noise, among others.
+Each corruption is characterized by 5 severity levels, yielding a total of 75 unique testing scenarios per dataset.
+Within each severity level, there are 10,000 images, contributing to a comprehensive evaluation of the model’s robustness against a variety of corruption types and intensities.
+Finally, we test our method on ImageNet-C to evaluate its performance on a larger dataset with 1,000 classes.
+Finally, we examine the VisDA-C dataset’s Peng et al.
+(2018) two domain shifts: simulated (3D) and video (YT).
+The former comprises a set of 152,397 images rendered in 3D across 12 different classes.
+The latter includes 72,372 YouTube video frames spanning the same categories.
+This dataset presents an important challenge, as it bridges the gap of the type of imagery that a model can be applied on.
+Appendix B Unsupervised clustering
 In Fig.
-3, tSNE visualizations of data points are shown.
+4, tSNE visualizations of data points are shown.
 We show how the distribution of data points change after adaptation, which improves the accuracy of class predictions and facilitates the assignment of ground truth labels.
@@
 Panels (c) and (d) demonstrate the actual labels in the absence of adaptation and following adaptation of the representations, respectively.
-Appendix B Evolution of similarity matrices
-A fundamental component of CLIPArTT is the employment of image-to-image and text-to-text similarity matrices to build new pseudo labels.
-In Fig.
-4, we show an example of a batch of five images.
-While the largest image-to-image similarity values concentrate on the main diagonal, the text-to-text similarity using the $K$ most confident classes for the new text prompt, highly impacts the correlation between images and text (CLIPArTT logits).
-Figure 4: Similarity matrices before (a) and after (b) adaptation using $K=5$ classes.
-There is an increase in the confidence of the logits when using a pseudolabel from the new prompt with the $K$ most confident classes.
-Appendix C Comprehensive experimental results
+Appendix C Additional settings
+We explore additional experimental settings to further explore the strengths and weaknesses of CLIPArTT.
+Although these scenarios deviate from the standard practices in the TTA literature, they are useful to evaluate a method’s performance in potentially challenging real-world applications.
+For most of the following experiments, we utilize the CIFAR-10 dataset’s variants, unless otherwise is mentioned.
+C.1 Imbalanced batch instances
+CLIPArTT adapts to batch data in a trandsductive manner by computing the image-to-image similarity.
+The diversity of the batches cannot be ensured due to their finite size.
+The opposite case naturally arises in large scale classification datasets such as ImageNet, where including at least one image per class in a 128--size batch is impossible.
+On the other hand, adaptation to highly imbalanced batches is understudied.
+In this experiment, we force extreme imbalances in the batches, by allowing only $C$ randomly chosen classes to be present in the batch.
+As CLIPArTT is dependant also on the most probable predictions, it is expected that the miss-classification due to the imbalance would drive the model to directions opposite to the actual correct classes.
+In this experiment, we evaluate with $C=\{2,3,4,5\}$ as the possible number of classes to be present in the batches, and compare CLIPArTT against TENT.
+Final results are shown in Table 9).
+It can be observed that TENT shows a higher robustness to class imbalance, as for each prediction, it depends less on the other images’ predictions (i.e., induction).
+Interestingly, we observe a trend where TENT’s performance decrease when more classes are present (e.g., $C=5$ instead of $C=2$), whereas CLIPArTT increase its performance as $C$ grows.
+CLIPArTT TENT
+2 3 4 5 2 3 4 5
+Original 76.99 81.12 84.37 86.92 97.07 96.14 95.27 94.09
+CIFAR-10.1 74.21 78.96 82.57 85.14 96.47 94.33 92.56 90.68
+Gaussian Noise 47.12 50.83 54.18 58.63 63.76 49.49 49.92 47.15
+Shot Noise 48.67 53.30 56.37 60.24 69.89 55.83 56.68 53.55
+Impulse Noise 45.98 49.82 53.15 57.48 66.87 54.47 54.77 52.88
+Defocus Blur 60.80 67.78 71.24 74.85 89.19 85.39 83.05 81.44
+Glass Blur 47.82 53.07 56.45 60.12 77.11 65.34 62.72 58.78
+Motion Blur 60.85 66.82 70.67 73.43 83.33 79.88 79.28 75.89
+Zoom Blur 56.78 66.52 70.43 72.94 86.88 84.15 81.89 81.35
+Snow 59.20 66.38 71.29 73.97 91.84 87.60 85.32 82.45
+Frost 60.02 68.08 72.12 75.15 92.50 88.62 86.18 84.25
+Fog 58.57 66.11 70.54 73.12 91.26 87.01 84.07 83.12
+Brightness 72.29 77.62 80.75 83.55 95.45 93.82 92.29 91.15
+Contrast 62.18 71.68 75.08 77.64 89.75 87.11 85.72 83.25
+Elastic Transform 50.85 58.91 62.12 65.12 88.17 82.05 78.13 76.05
+Pixelate 51.75 57.20 60.42 63.08 77.35 71.38 69.22 66.85
+JPEG Compression 49.63 56.08 59.32 62.05 79.98 72.34 69.85 68.17
+Table 9: Accuracy on imbalance datasets with different numbers of available classes at each batch.
+C.2 Open-Set classification
+Another interesting scenario is open-set classification, where images from classes that are not considered are present in the batch.
+These images could potentially affect performance, specially in transductive methods and particularly in CLIPArTT, as the image-to-image and text-to-text similarities are combined.
+The only available text prompts are wrongly assigned to the out-of-distribution (OOD) samples.
+To measure the effectiveness of our method in this scenario, we utilize SVHN-C, a corrupted version of the SVHN Netzer et al.
+(2011) dataset, analogous to CIFAR-10.C.
+For each batch of 128 image, 128 additional OOD images are included.
+Both parts are used at the same time for adaptation, and accuracy is only measured on the first half.
+As seen in Table 10, CLIPArTT defeats TENT’s open-set accuracy by a significant margin. This encouraging result suggests that our method is robust to semantic out-of-distribution perturbations.
+CLIPArTT TENT
+Gaussian Noise 59.72 $\pm$0.05 41.27 $\pm$0.03
+Shot Noise 62.11 $\pm$0.07 48.14 $\pm$0.05
+Impulse Noise 55.23 $\pm$0.10 47.86 $\pm$0.07
+Defocus Blur 75.44 $\pm$0.07 72.03 $\pm$0.05
+Glass Blur 60.12 $\pm$0.07 42.08 $\pm$0.08
+Motion Blur 74.60 $\pm$0.05 65.71 $\pm$0.05
+Zoom Blur 76.15 $\pm$0.05 71.45 $\pm$0.06
+Snow 74.69 $\pm$0.05 73.99 $\pm$0.08
+Frost 77.72 $\pm$0.05 74.49 $\pm$0.09
+Fog 72.87 $\pm$0.03 70.57 $\pm$0.04
+Brightness 85.70 $\pm$0.09 82.34 $\pm$0.05
+Contrast 76.49 $\pm$0.04 71.67 $\pm$0.04
+Elastic Transform 67.42 $\pm$0.01 65.99 $\pm$0.12
+Pixelate 62.02 $\pm$0.11 54.25 $\pm$0.08
+JPEG Compression 60.81 $\pm$0.07 54.09 $\pm$0.10
+Average 69.41 $\pm$8.86 62.40 $\pm$13.20
+Table 10: Open-set accuracy on CIFAR-10-C when using SVHN-C as the OOD dataset.
+C.3 Generalization after adaptation
+In TTA, the classification performance is evaluated directly on the adapted set of images in an episodic manner.
+However, a high accuracy on this set does not necessarily guarantee a good performance in a separate set of images unknown to the model.
+[…]
+While TENT obtains a higher accuracy on the adaptation splits of natural images (i.e., CIFAR-10 and CIFAR-10.1), following a similar trend as in the main experiments’ results, the difference in the generalization accuracy is smaller with respect to CLIPArTT’s.
+Moreover, our method demonstrates a consistently better generalization and adaptation accuracy on corrupted samples, representing stronger domain shifts.
+CLIPArTT TENT
+Acc. Acc. Gen. Acc. Acc. Gen.
+Original 95.43 $\pm$0.01 71.23 $\pm$0.0008 97.02 $\pm$0.02 72.68 $\pm$0.0023
+CIFAR-10.1 88.75 $\pm$0.85 65.87 $\pm$0.8083 90.49 $\pm$0.38 67.07 $\pm$1.1
+Gaussian Noise 63.79 $\pm$0.20 46.87 $\pm$0.0008 45.30 $\pm$0.05 30.83 $\pm$0.0030
+Shot Noise 66.38 $\pm$0.10 79.80 $\pm$0.0038 50.65 $\pm$0.18 35.85 $\pm$0.0025
+Impulse Noise 59.31 $\pm$0.19 44.56 $\pm$0.0023 51.91 $\pm$0.20 37.99 $\pm$0.0037
+Defocus Blur 81.26 $\pm$0.03 60.13 $\pm$0.0016 84.48 $\pm$0.07 60.32 $\pm$0.0028
+Glass Blur 65.38 $\pm$0.04 49.47 $\pm$0.0020 56.68 $\pm$0.26 41.32 $\pm$0.0035
+Motion Blur 80.41 $\pm$0.12 60.08 $\pm$0.0022 75.70 $\pm$0.13 55.72 $\pm$0.0013
+Zoom Blur 81.97 $\pm$0.05 60.63 $\pm$0.0009 81.14 $\pm$0.13 59.33 $\pm$0.0007
+Snow 82.06 $\pm$0.14 60.83 $\pm$0.0014 83.08 $\pm$0.13 61.08 $\pm$0.0008
+Frost 83.85 $\pm$0.19 61.73 $\pm$0.0009 84.70 $\pm$0.11 66.75 $\pm$0.0006
+Fog 90.04 $\pm$0.27 59.48 $\pm$0.0027 81.56 $\pm$0.09 60.69 $\pm$0.0033
+Brightness 91.73 $\pm$0.04 68.01 $\pm$0.0017 92.93 $\pm$0.03 68.80 $\pm$0.0018
+Contrast 82.63 $\pm$0.07 61.28 $\pm$0.0025 83.53 $\pm$0.06 62.97 $\pm$0.0009
+Elastic Transform 74.04 $\pm$0.22 54.87 $\pm$0.0011 74.08 $\pm$0.10 54.92 $\pm$0.0009
+Pixelate 70.58 $\pm$0.02 51.53 $\pm$0.0042 67.35 $\pm$0.18 50.19 $\pm$0.0004
+JPEG Compression 67.32 $\pm$0.06 49.56 $\pm$0.0023 66.25 $\pm$0.14 48.85 $\pm$0.0021
+Average 75.38$\pm$9.04 55.92 $\pm$6.62 71.78 $\pm$14.14 52.77 $\pm$11.10
+Table 11: Generalization results on CIFAR-10 variants.
+Each batch is divided into an adaptation split and a validation split.
+Accuracy (Acc.) is measured on the former after adaptation, whilst the generalization accuracy (Acc.
+Gen.) is measured on the later.
+Appendix D Computational Cost
+In this section, we compare the computational cost of CLIPArTT with other TTA methods through a thorough evaluation under consistent conditions, using an NVIDIA A6000 GPU within the same Python environment.
+The provided table 12 compares adaptation time, memory usage, and the number of learnable parameters across various TTA methods, including our proposed CLIPArTT.
+The results demonstrate that CLIPArTT maintains competitive adaptation time and memory usage relative to other approaches, such as TENT and TPT.
+Method Adaptation Time Memory Pct. of Learnable Parameters
+TENT 0.28 s 1.5 GB 0.026%
+TPT 0.26 s 1.7 GB 0.001%
+CLIPArTT 0.55 s 1.7 GB 0.026%
+Table 12: Comparison of Computational Cost.
+Appendix E Comprehensive experimental results
 We present comprehensive tables containing all the detailed information about results that was summarized in the main paper.